### v1

# RILe: Reinforced Imitation Learning

Abstract
Reinforcement Learning has achieved significant success in generating complex behavior but often requires extensive reward function engineering. Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator. Employing discriminators increases their data- and computational efficiency over the standard approaches; however, results in sensitivity to imperfections in expert data. We propose RILe, a teacher-student system that achieves both robustness to imperfect data and efficiency. In RILe, the student learns an action policy while the teacher dynamically adjusts a reward function based on the student’s performance and its alignment with expert demonstrations. By tailoring the reward function to both performance of the student and expert similarity, our system reduces dependence on the discriminator and, hence, increases robustness against data imperfections. Experiments show that RILe outperforms existing methods by 2x in settings with limited or noisy expert data.
1 Introduction
Reinforcement Learning (RL) offers a framework for learning behavior by maximizing a reward function. In recent years, deep reinforcement learning has demonstrated remarkable success in replicating complex behaviors, including playing Atari games, chess, and Go [1, 2]. However, designing a reward function is a tedious and challenging task, as predicting the policy outcome from a manually crafted reward function is difficult.
To address this issue, Imitation Learning (IL) leverages expert demonstrations for learning a policy. Since vast amounts of expert data are required to learn expert behaviors accurately, Adversarial Imitation Learning (AIL) approaches are proposed as a data-efficient alternative to IL [3]. AIL employs a discriminator to measure similarity between learned behavior and expert behavior, and rewards the agent in training accordingly. However, the discriminator tends to overfit to the dynamics of the demonstration setting, limiting its ability to generalize to new environments [4]. Furthermore, AIL is vulnerable to noise or imperfections in expert data, as the discriminator rewards the agent to perfectly mimic expert data, including its flaws.
Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering. In contrast to IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that leads to expert behavior. Thus, the reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior. This iterative process renders IRL computationally expensive [5]. Adversarial Inverse Reinforcement Learning (AIRL) [4] addresses this inefficiency by introducing a discriminator that allows for simultaneous learning of the policy and reward function. By implementing a special structure on the discriminator, AIRL can accurately recover the reward function of an expert, thereby improving generalization to environments with dynamics that vary from the demonstration setting, unlike AIL. However, in contrast to IRL, the reward function in AIRL is not learned independently but is directly derived from the output of the discriminator. Like AIL, AIRL suffers from the data-sensitivity issue associated with the discriminator, which limits its performance when expert data is imperfect.
To learn behaviors from imperfect data in a computationally efficient way, we propose Reinforced Imitation Learning (RILe). The primary motivation of RILe is to maintain computational efficiency and scalability through an adversarial framework, while learning a reward function from expert behavior (akin to IRL) as a separate mechanism. Our framework comprises two interacting agents: a student agent that learns a replicating policy and a teacher agent that learns a reward function and guides the student agent during training. For computational efficiency, we employ a discriminator that distinguishes expert data from student roll-outs as the reward function of the teacher. The core idea behind RILe is that the discriminator guides the reward learning process instead of directly guiding the action policy learning. Our contributions over AIL and AIRL are two-fold:
  1. 1.
    We learn the reward function (teacher agent) of the replicating policy (student agent) as an independent function in an adversarial IL/IRL setting. Using RL, the teacher agent learns to provide rewards that maximize the cumulative reward from the discriminator. The resulting reward function can provide meaningful guidance even when the discriminator overfits to noisy expert data, enabling the replicating policy to navigate and overcome sub-optimal states suggested by imperfect expert demonstrations.
  2. 2.
    The learned reward function in RILe adds an extra layer of flexibility to adversarial IL/IRL settings by allowing the teacher to customize rewards according to the developmental stage of the replicating agent, facilitating more accurate imitation of expert behavior (i.e., "online learned reward shaping").
Our experimental evaluation compares RILe to the state-of-the-art methods of AIL and AIRL, specifically GAIL [3] and AIRL [4], across three distinct sets of experiments: (1) imitating noisy expert data in a grid world, (2) imitating motion-capture data in continuous control tasks, and (3) learning to play image-based Atari games with discrete actions. Experimental results reveal that our approach outperforms baselines by 2x, especially when expert data is limited. We further demonstrate that RILe successfully learns expert-like behavior from noisy and misleading expert data, whereas the baseline methods fail to do so.
2 Related Work
We review the literature on learning expert behavior from demonstrations. Commonly, expert demonstrations are sourced either through direct queries to the expert in any observable state or by collecting sample trajectories demonstrated by the expert. We present related work that aligns with the most prevalent approaches of the latter setting, namely Imitation Learning and Inverse Reinforcement Learning. Both IL and IRL form the conceptual foundation of RILe.
Offline reinforcement learning also learns policies from data, which may include expert demonstrations. In contrast to our setting, its main goal is to learn a policy without any online interactions with the environment. We refer the reader to [6] for an overview of offline RL.
Imitation Learning
The earliest work on imitation learning introduced Behavioral Cloning (BC) [7], which aims to learn a policy congruent with expert demonstrations through supervised learning. DAgger [8] proposes the aggregation of expert demonstrations with policy experiences during the training of the policy for improving generalization over expert demonstrations. GAIL introduces and adversarial imitation learning method, where a discriminator aims to understand whether queried behavior stems from a policy or from expert demonstrations, while a generator tries to fool the discriminator by learning a policy that exhibits expert-like behavior [3]. InfoGAIL extends upon GAIL by extracting latent factors from expert behavior and employing them during imitation learning [9]. Deep Q-learning from Demonstrations (DQfD) proposes to first pre-train the learning agent using expert demonstrations, followed by a subsequent policy optimization through interactions with the environment [10]. ValueDice introduces an off-policy imitation learning method using a distribution-matching objective between policy and expert behavior [11].
Although the field of imitation learning has seen advancements, the need for high-quality expert data and data efficacy remain open challenges [5]. Moreover, the limited generalization capability of IL approaches persists [12]. We address these limitations related to IL’s data efficacy and data sensitivity by introducing an intermediary teacher agent that adapts itself based on the current status of the student agent, enabling the teacher to effectively guide the student beyond the space covered by the expert state-action pairs.
Inverse Reinforcement Learning
IRL is introduced in [13] to learn the intrinsic reward function of an expert and acquire the expert policy from it by iteratively optimizing an agents policy and then reward function. Apprenticeship learning builds on IRL and is proposed to represent the reward function as a linear combination of features [14]. Maximum Entropy Inverse Reinforcement Learning explicitly addresses the noise in expert demonstrations to better recover the reward function [15]. Several works extended IRL by incorporating negative demonstrations into the learning process to further improve behavior modeling [16, 17, 18]. Guided Cost Learning uses a neural network to approximate the reward function by using maximum entropy methods across continuous state-action spaces. [19]. AIRL proposes an adversarial reward learning framework to address the scalability issues of classical approaches [4]. IQ-Learn combines expert data with agent experiences to learn both reward and policy with a single Q-function [20]. Additionally, a pipeline is proposed that enables IRL to handle unstructured, real-world data [21]. XIRL introduces cross-embodiment scenarios, opening up a new direction in IRL research [22].
Despite the advancements in IRL, the computational efficacy of the learning process and scalability to complex problems remain open challenges [23]. The main reason for these limitations is the iterative sequential learning framework employed in IRL. We solve this efficacy problem by learning the policy and the reward function, via training a student agent and a teacher agent, in a single joint learning process.
3 Background
3.1 Preliminaries
A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$. $S$ is the state space that consists of all possible environment states $s$, and $A$ is action space that contains all possible environment actions $a$. $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function. $T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$. $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor. The policy $\pi=\pi(a|s):S\ \rightarrow\ A$ is a mapping from states to actions. In this work, we consider $\gamma$-discounted infinite horizon settings. Following [3], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $E_{\pi}[R(s,a)]\triangleq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$. The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$. We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [19].
Our work considers an imitation learning problem from expert trajectories, each consisting of states $s$ and actions $a$. The set of expert trajectories $\tau_{E}$ are sampled from an expert policy $\pi_{E}\in\Pi$, where $\Pi$ is the set of all possible policies. We assume that we have access to $m$ expert trajectories, all of which have $n$ time-steps, $\tau_{E}=\{(s^{i}_{0},a^{i}_{0}),(s^{i}_{1},a^{i}_{1}),\ldots,(s^{i}_{n},a^{i}_{n})\}_{i=1}^{m}$.
3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$. In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 24]. The RL problem with a parameterized reward function and entropy regularization is defined as
$\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{argmax}_{\pi}E_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ of an optimal expert policy $\pi_{E}$, inverse reinforcement learning $\text{IRL}(\tau_{E})$ seeks to recover a reward function $R_{\theta}^{*}(s,a)$ that maximally rewards behavior that was sampled from the expert policy $\pi_{E}$. In other words, IRL aims to find a reward function $R_{\theta}^{*}(s,a)$ that satisfies
$E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ (2)
When IRL recovers the optimal reward function, $R_{\theta}^{*}(s,a)$ would result in a policy that exhibits expert behavior when optimized with reinforcement learning: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{argmax}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})].$
Since only the trajectories created by the expert policy are observed, expectations are estimated from the samples in $\tau_{E}$. In other words, IRL seeks to find the reward function in which the expert policy performs better than any other policy, where the expectation of the expert is estimated through samples. With entropy regularization $H(\pi)$, maximum casual entropy inverse reinforcement learning [15] can be defined as
$\text{IRL}(\tau_{E})=\operatorname*{argmax}_{R_{\theta}(s,a)\in\mathbb{R}}\left(E_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(E_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (3)
3.4 Adversarial Imitation Learning (AIL)
In contrast to inverse reinforcement learning, imitation learning aims to directly approximate the expert policy from given expert trajectory samples. It can be formulated as
$\text{IL}(\tau_{E})=\operatorname*{argmin}_{\pi}E_{(s,a)\sim\tau_{E}}[L(\pi(\cdot|s),a)],$ (4)
where $L$ is a loss function, that captures the difference between policy and expert data.
GAIL [3] extends imitation learning to an adversarial setting by quantifying the difference between policies of the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$. The discriminator’s function is to differentiate between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$. Its optimization problem is defined as
$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\notin\tau_{E}}[\text{log}(1-D_{\phi}(s,a))].$ (5)
Using the discriminator, the goal of GAIL is to find the optimal policy that minimizes this difference metric while maximizing an entropy constraint by training the discriminator and the policy at the same time. The optimization problem can be formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$, represented by
$\min_{\pi}\max_{\phi}E_{\pi}[\text{log}\ D_{\phi}(s,a)]+E_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (6)
In other words, the reward function that is maximized by the policy is defined as a similarity function, expressed as $R(s,a)=-\text{log}\ (D_{\phi}(s,a))$.
4 RILe: Reinforced Imitation Learning
We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process. Our framework consists of three key components: a student agent, a discriminator, and a teacher agent (Figure 1).
Intuitively, the student agent learns the action policy and the teacher agent learns the reward function. Both agents are trained simultaneously via RL with the help of an adversarial discriminator. The discriminator enables training the policy and reward function in one learning process and permits replicating expert behavior through similarity rather than supervised state action mapping. However, uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning. This setup enables the teacher agent to learn a reward function separately by continuously adapting to changes in the student’s policy and guiding the student based on the discriminator’s feedback. As a result, the teacher can understand different regions of the state-action space through student’s interactions, effectively guiding it beyond the potentially noisy and imperfect expert state-action pairs.
In the following, we define the components of RILe and explain how they can efficiently learn behavior from imperfect data.
Figure 1: Framework overview. The framework consists of three key components: a student agent, a teacher agent, and a discriminator. (1) The student agent learns a policy $\pi_{S}$ by interacting with an environment. (2) Based on the state action pair of the student agent $s^{T}=(s^{S},a^{S})$, the teacher agent learns a reward function as a policy $\pi_{T}$ by (3) choosing an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$. (4) The teacher agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator. (5) The discriminator compares student state-action pair with expert demonstrations $s^{D}$ and (6) the rewards the teacher based on their similarity $r^{T}=D(s^{T})$.
Student Agent
The student agent aims to learn a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP. For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$. However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the teacher agent $\pi_{T}$. Therefore, the reward function is represented by the teacher policy. Thus, the student agent is guided by the actions of the teacher agent, i.e., the action of the teacher is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$. The optimization problem of the student agent is then defined as
$\min_{\pi_{S}}-E_{(s^{S},a^{S})\sim\pi_{S}}[\pi_{T}\left((s^{S},a^{S})\right)].$ (7)
Discriminator
The discriminator differentiates between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and state-action pairs stemming from the policy of the student $(s,a)\sim\pi_{S}$. In RILe, the discriminator is defined as a feed-forward deep neural network, parameterized by $\phi$. Hence, the optimization problem is
$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\sim\pi_{S}}[\text{log}(1-D_{\phi}(s,a))].$ (8)
To provide effective guidance, the discriminator needs to accurately distinguish whether a given state-action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$. The feasibility of this discrimination has been demonstrated by GAIL [3]. The according lemma and proof are presented in the Appendix A.
Teacher Agent
The teacher agent aims to guide the student to approximate expert behavior by operating as its reward mechanism. Since the teacher agent takes the role of a reward function for the student, a new MDP is defined for the teacher agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1. $A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value. $R_{T}(s,a)$ is the reward function where $s\in S_{T}$ and $a\in A_{T}$. $T_{T}=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as the state distribution upon taking action $a\in A_{T}$ in state $s\in S_{T}$. $K_{T}$ is the initial state distribution.
The teacher agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$. Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the teacher comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$. It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1. To help the teacher understand how its actions impact the reward it receives, we define the reward function such that it multiplies the discriminator’s output by a sigmoid function dependent on the teacher’s actions. Therefore, the teacher agent’s reward function is defined as $R^{T}=\text{log}(D_{\phi}(s^{T}))\sigma(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\sigma(a^{T})$ is the sigmoid function applied to the teacher’s own actions. Considering only the discriminator’s output in the teacher agent’s reward function would result in a reward independent of the teacher’s actions.
The optimization problem of the teacher can be defined as
$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\text{log}(D_{\phi}(s,a))\sigma(a^{T})].$ (9)
RILe
RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$. In RILe, the student policy $\pi_{S}$ and the teacher policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method. The training algorithm is given in Appendix C. Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as
$\pi_{S}^{*}=\operatorname*{argmax}_{\pi_{S}}E_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)\right].$ (10)
At the same time, the teacher agent aims to recover the optimal policy $\pi_{T}^{*}$ as
$\pi_{T}^{*}=\operatorname*{argmax}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\text{log}(D_{\phi}(s^{T}_{t}))\sigma(a^{T})\right].$ (11)
To prove that the student agent can learn expert-like behavior, we need to show that the teacher agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
Lemma 1: Given the discriminator $D_{\phi}$, the teacher agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
Proof for Lemma 1 The assumptions are presented in Appendix A. The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$. The role of the teacher is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student. The teacher’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
The Value and Q functions of the teacher, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards. The value function for the teacher’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
$V^{\theta_{T}}(s_{t})=\mathbb{E}_{\pi^{\theta_{T}}}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}\right],$ (12)
where $r_{t+k}=D_{\phi}(s_{t+k})$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
$Q^{\theta_{T}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{T}}}\left[V^{\theta_{T}}(s_{t+1})\mid s_{t},a_{t}\right].$ (13)
The teacher’s policy optimization is done by maximizing the following clipped surrogate objective function:
$\begin{split}L^{CLIP}(\theta_{T})=&\mathbb{E}_{(s_{t},a_{t})\sim\pi^{\theta_{T}}}\Big{[}\min\Big{(}\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})}A^{\theta_{T_{old}}}(s_{t},a_{t}),\\ &\qquad\qquad\qquad\text{clip}(\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon)A^{\theta_{T_{old}}}(s_{t},a_{t})\Big{)}\Big{]},\end{split}$ (14)
where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the teacher’s old policy parameters $\theta_{T_{old}}$.
By expressing the advantage using the reward from the discriminator, we explicitly tie the policy gradient updates to the discriminator’s output, emphasizing the shaping of $\pi^{\theta_{S}}$ to match $p_{\text{expert}}(s,a)$:
$A^{\theta_{T_{old}}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}(s_{t+1})\right]-V^{\theta_{T_{old}}}(s_{t}).$ (15)
$A^{\theta_{T_{old}}}((s^{S}_{t},a^{S}_{t}),a_{t})=D_{\phi}((s^{S}_{t},a^{S}_{t}))+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}((s^{S}_{t+1},a^{S}_{t+1}))\right]-V^{\theta_{T_{old}}}(s^{S}_{t},a^{S}_{t}).$ (16)
During each policy update, the objective in Equation 14 is maximized, driving parameter updates to favor actions that elicit higher rewards from the discriminator – effectively the actions that better align with expert behavior:
$\theta_{T}\leftarrow\theta_{T}+\alpha_{\pi}\nabla_{\theta_{T}}L^{CLIP}(\theta_{T}).$ (17)
The update rule in Equation 17, driven by cumulative rewards from the discriminator, incrementally adapts the teacher’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$. The teacher’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
5 Experiments
We evaluate the performance of RILe by addressing three key questions. First, can RILe recover a feasible policy when trained to imitate from noisy expert data? To answer this question, we compare RILe’s performance with common imitation learning baselines in a grid world setting where policies are learned from noisy expert data. Second, can RILe use expert-data explicitly to imitate expert behavior? We use MuJoCo [25, 26] to evaluate RILe’s performance when expert data is leaked to the agents. Third, is RILe efficient and scalable to high-dimensional continuous control tasks? We answer this by using LocoMujoco [27] as the benchmark for evaluating RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks.
Baselines We compared RILe with four baseline methods: Behavioral cloning (BC [7, 28]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [4]) and inverse reinforcement learning (IQ-Learn [20]).
5.1 Noisy Expert Data
(a) Expert traj.
(b) RILe traj.
(c) GAIL traj.
(d) AIRL traj.
(e) IQLearn traj.
(f) RILe val.
(g) GAIL val.
(h) AIRL val.
(i) IQLearn val.
Figure 2: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death. (c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish. (b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal. Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively. The optimal path, derived from the reward of the teacher or discriminator, is highlighted with green lines.
To demonstrate the advantage of using RL to learn the reward function in RILe, as opposed to deriving the reward directly from the discriminator in AIL and AIRL, we designed a 5x5 MiniGrid experiment. The grid consists of 4 lava tiles that immediately kill the agent if it steps in it, representing terminal conditions. The goal condition of the environment is reaching the green tile.
The expert demonstrations are imperfect, depicting an expert that passes through a lava tile without being killed and still reaches the green goal tile. Using this data, we trained the adversarial approaches with a perfect discriminator, which provides a reward of $0.99$ if the visited state-action pair stems from the expert and $0.01$ otherwise. These values were chosen over 1 and 0 because both AIRL and GAIL use the logarithm of the discriminator output to calculate rewards.
Results are presented in Fig. 2. The value graphs (Fig. 2e-g) are attained by computing the value of each grid cell $c_{i}$ as $\sum_{a\in A}D(c_{i},a)$ for AIRL and GAIL, and $\sum_{a\in A}\pi_{T}(c_{i},a)$ for RILe. Fig. 2a shows the expert trajectory.
GAIL (Fig. 2c), AIRL (Fig. 2d) and IQLearn (Fig. 2e) fail to reach the goal, as their agents either become stuck or are directed into lava. In contrast, RILe (Fig. 2d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data. The difference in the value graphs between RILe and the baselines intuitively explains this outcome. In AIL and AIRL (Fig. 2f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly. Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory. In contrast, RILe’s teacher agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting. By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig. 2f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal. Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration. Since the student agent is guided by the teacher to also match the expert trajectory, it remains close to this path after passing the lava tiles.
5.2 Expert Data Leakage in Training
We investigate whether RILe can use expert data explicitly to imitate expert behavior better. Thus, we analyze how the performance and convergence time will be effected when we leak expert data to the replay buffers of both student and teacher agent. For this experiment, we employed MuJoCo’s Humanoid environment [26, 25] and used a single trajectory from the expert data of [20]. The expert data is leaked to agents’ replay buffers with following percentages: $25\%,50\%,75\%,100\%$. $25\%$ signifies that $25\%$ of the replay buffer’s data is sourced from the expert trajectory, while the remaining $75\%$ originates from the agents’ interactions with the environment. $100\%$ indicates that the agents’ environment interactions are entirely disregarded, with only expert data being used for training.
As shown in Figure 3, leaking expert data to buffers slightly worsened performance but significantly reduced the convergence time. However, when environment interactions were entirely replaced with expert data, the policy’s performance declined substantially. For comparison, results from IQLearn and Behavioral Cloning are also presented, as both methods rely exclusively on expert data for training. Neither method, however, achieved the same level of performance as RILe.
Normalized Scores
Normalized Steps
Figure 3: Scores (left) representing the total rewards achieved, and convergence steps (right) representing the number of steps to convergence, are presented against the percentage of leaked expert data. For RILe, the x-axis shows the percentage of leaked expert data in the student agent’s buffer, while the y-axis shows it in the teacher agent’s buffer. Scores/Steps are normalized by no-leakage training results of RILe. Behavioral Cloning fails to achieve a comparable score to RILe. While IQLearn converges faster than the no-leakage RILe, it still falls short in terms of overall performance.
5.3 Motion-Capture Data Imitation for Robotic Continuous Control
In LocoMujoco, the objective is to learn policies that can accurately imitate the collected MoCap data of various robotic tasks. LocoMujoco is challenging due its high dimensionality. The MoCap data only consists of states of the robot, not its actions, which prevents the use of Behavioral Cloning and IQLearn in this analysis, since they both requires to have full expert data with states and actions. It should be noted that LocoMujoco environments do not have terminal or goal conditions.
Results are presented in Table 1 for both test seeds and validation seeds. RILe outperforms baselines especially for new initial conditions introduced by varying seeds.
Table 1: Test and validation results on seven LocoMujoco tasks.
Test Seeds Validation Seeds
RILe GAIL AIRL RILe GAIL AIRL Expert
Atlas Walk 870.6 792.7 300.5 895.4 918.6 356 1000
Atlas Carry 850.8 669.3 256.4 889.7 974.2 271.9 1000
Talos Walk 842.5 442.3 102.1 884.7 675.5 103.4 1000
Talos Carry 220.1 186.3 134.2 503.3 338.5 74.1 1000
UnitreeH1 Walk 898.3 950.2 568.1 980.7 965.1 716.2 1000
UnitreeH1 Carry 788.3 634.6 130.5 850.6 637.4 140.9 1000
Humanoid Walk 831.3 181.4 80.1 970.3 216.2 78.2 1000
6 Discussion
We have demonstrated through our experiments that RILe outperforms baseline models across various settings. While RILe surpasses the state-of-the-art adversarial IL/IRL methods, it also has its limitations. The main challenge is training the policy with a changing reward function, which leads to unstable policy updates for the student agent. To address this issue, we formulate the teacher’s reward by correlating its actions with the output of the discriminator, instead of only relying on the discriminator. However, instability remains a challenge and causes slower learning of the action policy in RILe compared to AIL. Future work should establish bounds for the reward agent updates to enhance the stability of the student agent’s learning process.
Moreover, balancing the learning rates of the discriminator and the policies is challenging, particularly because the discriminator tends to overfit the data quickly. This overfitting makes it difficult for the teacher agent to find a reward for the student that can challenge the discriminator. While adjusting the discriminator’s learning rate on a case-by-case basis helps, future research should focus on a fundamental solution, such as replacing the discriminator with a more robust similarity component.
Uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning. Therefore, the reward function is learned, not directly derived from the discriminator. This distinction allows RILe to provide meaningful guidance even when the discriminator overfits to noisy expert data. In the LocoMujoco benchmark, where perfect expert data is available, we argue that RILe’s ability to account for the agent’s exploration efforts during training, in addition to its behavioral similarity with the expert, results in a superior generalization capability compared to AIL and AIRL.
Appendix A Justification of RILe
Assumptions:
  • The discriminator loss curve is complex and the discriminator function, $D_{\phi}(s,a)$, is sufficiently expressive since it is parameterized by a neural network with adequate capacity.
  • For the teacher’s and student’s policy functions $(\pi^{\theta_{T}})$ and $(\pi^{\theta_{S}})$, and the Q-functions $(Q^{\theta_{S}})$, each is Lipschitz continuous with respect to its parameters with constants $(L_{\theta_{T}}),(L_{\theta_{S}}),and(L_{Q})$, respectively. This means for all $(s,a)$ and for any pair of parameter settings $(\theta,\theta^{\prime}):[|\pi^{\theta}(s,a)-\pi^{\theta^{\prime}}(s,a)|\leq L_{\theta}|\theta-\theta^{\prime}|,][|Q^{\theta}(s,a)-Q^{\theta^{\prime}}(s,a)|\leq L_{Q}|\theta-\theta^{\prime}|.]$
A.1 Lemma 2:
The discriminator $D_{\phi}$, parameterized by $\phi$ will converge to a function that estimates the probability of a state-action pair being generated by the expert policy, when trained on samples generated by both a student policy $\pi^{\theta_{S}}$ and an expert policy $\pi_{E}$.
Proof for Lemma 2: The training objective for the discriminator is framed as a binary classification problem over mini-batches from expert demonstrations and student-generated trajectories. The discriminator’s loss function $\mathcal{L}_{D}(\phi)$ is the binary cross-entropy loss, which for a mini-batch of size $n$ is defined as:
$\mathcal{L}_{D}(\phi)=-\frac{1}{n}\sum_{i=1}^{n}\Big{[}y_{i}\log(D_{\phi}(s_{i},a_{i}))+(1-y_{i})\log(1-D_{\phi}(s_{i},a_{i}))\Big{]},$ (18)
where $(s_{i},a_{i})$ are sampled state-action pairs from the combined replay buffer $\mathcal{D}=\mathcal{D}_{S}\cup\mathcal{D}_{E}$, with corresponding labels $y_{i}$ indicating whether the pair is from the expert $(y_{i}=1)$ or the student $(y_{i}=0)$. The stochastic gradient descent update rule for optimizing $\mathcal{L}_{D}(\phi)$ is then given by:
$\phi\leftarrow\phi-\eta_{D}\nabla_{\phi}\mathcal{L}_{D}(\phi),$ (19)
where $\eta_{D}$ is the learning rate for the discriminator.
Through iterative updates, $D_{\phi}(s,a)$ will converge to $P(\pi_{E}|s,a)$, provided the minimization of $\mathcal{L}_{D}(\phi)$ progresses according to the theoretical foundations of stochastic gradient descent. The convergence relies on the assumption that the student’s policy $\pi^{\theta_{S}}$ and the expert policy $\pi_{E}$ induce stationary distributions of state-action pairs, such that the discriminator’s data source is consistently representing both policies over time.
By minimizing $\mathcal{L}_{D}(\phi)$, we seek $\phi^{*}$ such that:
$\phi^{*}=\arg\min_{\phi}\mathcal{L}_{D}(\phi).$ (20)
Under typical conditions for convergence in stochastic gradient descent, the convergence to a local minimum or saddle point can be guaranteed. The discriminator’s ability to distinguish between student and expert pairs improves as $\mathcal{L}_{D}(\phi)$ is minimized, implying that $\lim_{n_{batch}\rightarrow\infty}D{\phi^{*}}(s,a)=P(\pi_{E}|s,a)$, where $n_{batch}$ is the number of batches.
Appendix B Compute Resources
For the training of RILe and baselines, following computational sources are employed:
  • AMD EPYC 7742 64-Core Processor
  • 1 x Nvidia RTX6000 GPU
  • 32GB Memory
Appendix C Algorithm
Algorithm 1 RILe Training Process
1:Initialize student policy $\pi_{S}$ and teacher policy $\pi_{T}$ with random weights, and the discriminator $D$ with random weights.
2:Initialize an empty replay buffer $B$
3:for each iteration do
4:     Sample trajectory $\tau_{S}$ using current student policy $\pi_{S}$
5:     Store $\tau_{S}$ in replay buffer $B$
6:     for each transition $(s,a)$ in $\tau_{S}$ do
7:         Calculate student reward $R^{S}$ using teacher policy:
$R^{S}\leftarrow\pi_{T}$ (21)
8:         Update $\pi_{S}$ using policy gradient with reward $R^{S}$
9:     end for
10:     Sample a batch of transitions from $B$
11:     Train discriminator $D$ to classify student and expert transitions
$\max_{D}E_{\pi_{S}}[\log(D(s,a))]+E_{\pi_{E}}[\log(1-D(s,a))]$ (22)
12:     for each transition $(s,a)$ in $\tau_{S}$ do
13:         Calculate teacher reward $R^{T}$ using discriminator:
$R^{T}\leftarrow\log(D(s,a))$ (23)
14:         Update $\pi_{T}$ using policy gradient with reward $R^{T}$
15:     end for
16:end for

### v2

@@
 Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator.
-Employing discriminators increases their data- and computational efficiency over the standard approaches; however, results in sensitivity to imperfections in expert data.
-We propose RILe, a teacher-student system that achieves both robustness to imperfect data and efficiency.
-In RILe, the student learns an action policy while the teacher dynamically adjusts a reward function based on the student’s performance and its alignment with expert demonstrations.
-By tailoring the reward function to both performance of the student and expert similarity, our system reduces dependence on the discriminator and, hence, increases robustness against data imperfections.
-Experiments show that RILe outperforms existing methods by 2x in settings with limited or noisy expert data.
+However, these methods struggle in complex tasks where randomly sampling expert-like behaviors is challenging.
+This limitation stems from their reliance on policy-agnostic discriminators, which provide insufficient guidance for agent improvement, especially as task complexity increases and expert behavior becomes more distinct.
+We introduce RILe (Reinforced Imitation Learning environment), a novel trainer-student system that learns a dynamic reward function based on the student’s performance and alignment with expert demonstrations.
+In RILe, the student learns an action policy while the trainer, using reinforcement learning, continuously updates itself via the discriminator’s feedback to optimize the alignment between the student and the expert.
+The trainer optimizes for long-term cumulative rewards from the discriminator, enabling it to provide nuanced feedback that accounts for the complexity of the task and the student’s current capabilities.
+This approach allows for greater exploration of agent actions by providing graduated feedback rather than binary expert/non-expert classifications.
+By reducing dependence on policy-agnostic discriminators, RILe enables better performance in complex settings where traditional methods falter, outperforming existing methods by 2x in complex simulated robot-locomotion tasks.
 1 Introduction
-Reinforcement Learning (RL) offers a framework for learning behavior by maximizing a reward function.
-In recent years, deep reinforcement learning has demonstrated remarkable success in replicating complex behaviors, including playing Atari games, chess, and Go [1, 2].
-However, designing a reward function is a tedious and challenging task, as predicting the policy outcome from a manually crafted reward function is difficult.
-To address this issue, Imitation Learning (IL) leverages expert demonstrations for learning a policy.
-Since vast amounts of expert data are required to learn expert behaviors accurately, Adversarial Imitation Learning (AIL) approaches are proposed as a data-efficient alternative to IL [3].
-AIL employs a discriminator to measure similarity between learned behavior and expert behavior, and rewards the agent in training accordingly.
-However, the discriminator tends to overfit to the dynamics of the demonstration setting, limiting its ability to generalize to new environments [4].
-Furthermore, AIL is vulnerable to noise or imperfections in expert data, as the discriminator rewards the agent to perfectly mimic expert data, including its flaws.
+Reinforcement Learning (RL) has emerged as a powerful framework for teaching agents to perform complex tasks.
+In recent years, deep reinforcement learning has demonstrated remarkable success in replicating sophisticated behaviors, including playing Atari games, chess, and Go [1, 2].
+However, these achievements often come at a cost: the tedious and challenging process of designing reward functions, as predicting the policy outcome from a manually crafted reward function remains notoriously difficult.
+To overcome the reward engineering problem, Imitation Learning (IL) leverages expert demonstrations to learn a policy.
+Since vast amounts of expert data are required to accurately learn expert behaviors, Adversarial Imitation Learning (AIL) approaches, such as GAIL [3], have been proposed as data-efficient alternatives.
+AIL employs a discriminator to measure similarity between learned behavior and expert behavior, rewarding the agent accordingly.
+While computationally efficient, AIL methods suffer from a critical limitation: the policy-agnostic nature of their discriminators.
+The discriminator lacks any inherent incentive to guide the agent towards expert-like behavior, in contrast to engineered reward functions in RL.
+Consequently, AIL methods face challenges in complex tasks requiring extensive exploration to find optimal actions.
+For instance, in digital locomotion tasks, AIL methods often struggle to consistently replicate expert performance [4].
 Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering.
-In contrast to IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that leads to expert behavior.
-Thus, the reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior.
+Unlike IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that motivates the agent to acquire expert behaviors.
+The reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior.
 This iterative process renders IRL computationally expensive [5].
-Adversarial Inverse Reinforcement Learning (AIRL) [4] addresses this inefficiency by introducing a discriminator that allows for simultaneous learning of the policy and reward function.
-By implementing a special structure on the discriminator, AIRL can accurately recover the reward function of an expert, thereby improving generalization to environments with dynamics that vary from the demonstration setting, unlike AIL.
-However, in contrast to IRL, the reward function in AIRL is not learned independently but is directly derived from the output of the discriminator.
-Like AIL, AIRL suffers from the data-sensitivity issue associated with the discriminator, which limits its performance when expert data is imperfect.
-To learn behaviors from imperfect data in a computationally efficient way, we propose Reinforced Imitation Learning (RILe).
-The primary motivation of RILe is to maintain computational efficiency and scalability through an adversarial framework, while learning a reward function from expert behavior (akin to IRL) as a separate mechanism.
-Our framework comprises two interacting agents: a student agent that learns a replicating policy and a teacher agent that learns a reward function and guides the student agent during training.
-For computational efficiency, we employ a discriminator that distinguishes expert data from student roll-outs as the reward function of the teacher.
-The core idea behind RILe is that the discriminator guides the reward learning process instead of directly guiding the action policy learning.
-Our contributions over AIL and AIRL are two-fold:
+Adversarial Inverse Reinforcement Learning (AIRL) [6] attempts to address this inefficiency by introducing a discriminator that enables simultaneous learning of the policy and reward function.
+However, in AIRL, the reward function is tightly coupled to the discriminator, potentially limiting its ability to capture complex task structures or long-term dependencies and inheriting the limitations of a policy-agnostic discriminators.
+This highlights the need for a method that can learn a more flexible reward function without the computational overhead of traditional IRL methods.
+To overcome these challenges and effectively learn behaviors in complex settings, we propose Reinforced Imitation Learning (RILe) (Fig.
+1-(d)).
+RILe aims to combine the ability to learn a reward function that actively guides the agent to imitate expert behavior with the computational efficiency of adversarial frameworks.
+At the core of RILe is a novel trainer-student system designed to address the shortcomings of existing methods:
+  • A student agent that learns to replicate the expert’s policy via RL in the environment
+  • A trainer agent that learns a reward function via RL and guides the student agent during training
+By integrating the trainer-student dynamic, RILe decouples reward learning from policy learning and the discriminator, allowing each component to specialize and thereby overcome the limitations of policy-agnostic discriminators.
+While RILe utilizes a discriminator similar to those in adversarial frameworks, its role is fundamentally redefined.
+In RILe, the discriminator’s primary function is to provide feedback to the trainer agent by distinguishing expert data from student roll-outs.
+This feedback serves as the reward signal for the trainer, not directly influencing the student agent.
+The trainer leverages the discriminator’s feedback to learn a reward function that effectively guides the student agent.
+This approach enables more nuanced reward shaping, particularly beneficial in tasks requiring complex decision-making and extensive exploration.
+Our contributions are two-fold:
   1. 1.
-    We learn the reward function (teacher agent) of the replicating policy (student agent) as an independent function in an adversarial IL/IRL setting.
-Using RL, the teacher agent learns to provide rewards that maximize the cumulative reward from the discriminator.
-The resulting reward function can provide meaningful guidance even when the discriminator overfits to noisy expert data, enabling the replicating policy to navigate and overcome sub-optimal states suggested by imperfect expert demonstrations.
+    Decoupled Reward-function Learning: We introduce a novel approach where the trainer agent learns the reward function independently from both the student agent and the discriminator.
+Unlike existing methods that derive rewards directly from discriminator outputs, our trainer agent uses reinforcement learning to optimize the reward function based on the feedback from the discriminator.
+By focusing on long-term reward maximization, RL enables the trainer to distill inconsistent feedback from the discriminator into meaningful rewards, leading to better student performance.
   2. 2.
-    The learned reward function in RILe adds an extra layer of flexibility to adversarial IL/IRL settings by allowing the teacher to customize rewards according to the developmental stage of the replicating agent, facilitating more accurate imitation of expert behavior (i.e., "online learned reward shaping").
-Our experimental evaluation compares RILe to the state-of-the-art methods of AIL and AIRL, specifically GAIL [3] and AIRL [4], across three distinct sets of experiments: (1) imitating noisy expert data in a grid world, (2) imitating motion-capture data in continuous control tasks, and (3) learning to play image-based Atari games with discrete actions.
-Experimental results reveal that our approach outperforms baselines by 2x, especially when expert data is limited.
-We further demonstrate that RILe successfully learns expert-like behavior from noisy and misleading expert data, whereas the baseline methods fail to do so.
+    Dynamic Reward Customization: Our trainer agent dynamically adjusts rewards based on the student agent’s progress, facilitating a better learning experience and enabling accurate imitation of expert behavior in complex settings.
+This adaptive approach allows for more gradual learning, particularly in tasks where the optimal behavior may change depending on the agent’s current capabilities.
+We evaluate RILe against state-of-the-art methods in AIL, and AIRL, specifically GAIL [3], and AIRL [6].
+Our experiments span three distinct scenarios: (1) Tailoring a reward function dynamically in a discrete maze task, (2) Investigating the impact of expert data on the trainer-student dynamics in a humanoid locomotion task, and (3) Imitating motion-capture data in continuous control tasks.
+The results demonstrate RILe’s superior performance, especially in complex tasks, and its ability to learn an effective dynamic reward function where baseline methods fail.
 2 Related Work
-We review the literature on learning expert behavior from demonstrations.
-Commonly, expert demonstrations are sourced either through direct queries to the expert in any observable state or by collecting sample trajectories demonstrated by the expert.
-We present related work that aligns with the most prevalent approaches of the latter setting, namely Imitation Learning and Inverse Reinforcement Learning.
-Both IL and IRL form the conceptual foundation of RILe.
-Offline reinforcement learning also learns policies from data, which may include expert demonstrations.
-In contrast to our setting, its main goal is to learn a policy without any online interactions with the environment.
-We refer the reader to [6] for an overview of offline RL.
+We review literature on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), which form the conceptual foundation of RILe.
 Imitation Learning
-The earliest work on imitation learning introduced Behavioral Cloning (BC) [7], which aims to learn a policy congruent with expert demonstrations through supervised learning.
-DAgger [8] proposes the aggregation of expert demonstrations with policy experiences during the training of the policy for improving generalization over expert demonstrations.
-GAIL introduces and adversarial imitation learning method, where a discriminator aims to understand whether queried behavior stems from a policy or from expert demonstrations, while a generator tries to fool the discriminator by learning a policy that exhibits expert-like behavior [3].
-InfoGAIL extends upon GAIL by extracting latent factors from expert behavior and employing them during imitation learning [9].
-Deep Q-learning from Demonstrations (DQfD) proposes to first pre-train the learning agent using expert demonstrations, followed by a subsequent policy optimization through interactions with the environment [10].
-ValueDice introduces an off-policy imitation learning method using a distribution-matching objective between policy and expert behavior [11].
-Although the field of imitation learning has seen advancements, the need for high-quality expert data and data efficacy remain open challenges [5].
-Moreover, the limited generalization capability of IL approaches persists [12].
-We address these limitations related to IL’s data efficacy and data sensitivity by introducing an intermediary teacher agent that adapts itself based on the current status of the student agent, enabling the teacher to effectively guide the student beyond the space covered by the expert state-action pairs.
+Early work introduced Behavioral Cloning (BC) [7], which learns a policy congruent with expert demonstrations through supervised learning.
+DAgger [8] introduces data aggregation.
+GAIL [3] introduces adversarial methods, where a discriminator aims to discriminate expert demonstrations, while a generator tries to fool the discriminator.
+DQfD [9] proposes two-stage approach with pre-training, and ValueDice [10] uses a distribution-matching objective between policy and expert.
+Despite progress, IL faces challenges in efficacy and generalization [5, 11].
+RILe addresses these by introducing an adaptive teacher agent to guide the student beyond expert demonstrations.
 Inverse Reinforcement Learning
-IRL is introduced in [13] to learn the intrinsic reward function of an expert and acquire the expert policy from it by iteratively optimizing an agents policy and then reward function.
-Apprenticeship learning builds on IRL and is proposed to represent the reward function as a linear combination of features [14].
-Maximum Entropy Inverse Reinforcement Learning explicitly addresses the noise in expert demonstrations to better recover the reward function [15].
-Several works extended IRL by incorporating negative demonstrations into the learning process to further improve behavior modeling [16, 17, 18].
-Guided Cost Learning uses a neural network to approximate the reward function by using maximum entropy methods across continuous state-action spaces.
-[19].
-AIRL proposes an adversarial reward learning framework to address the scalability issues of classical approaches [4].
-IQ-Learn combines expert data with agent experiences to learn both reward and policy with a single Q-function [20].
-Additionally, a pipeline is proposed that enables IRL to handle unstructured, real-world data [21].
-XIRL introduces cross-embodiment scenarios, opening up a new direction in IRL research [22].
-Despite the advancements in IRL, the computational efficacy of the learning process and scalability to complex problems remain open challenges [23].
-The main reason for these limitations is the iterative sequential learning framework employed in IRL.
-We solve this efficacy problem by learning the policy and the reward function, via training a student agent and a teacher agent, in a single joint learning process.
+IRL, introduced by [12], learns the expert’s intrinsic reward function.
+Key developments include Apprenticeship Learning [13], Maximum Entropy IRL [14], and adversarial approaches like AIRL [6].
+Recent work explores handling unstructured data [15] and cross-embodiment scenarios [16].
+Despite advancements, IRL faces challenges in computational efficiency and scalability [17].
+RILe addresses these by jointly learning policy and reward function in a single process.
 3 Background
-3.1 Preliminaries
+3.1 Markov Decision Process
 A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$.
@@
 3.2 Reinforcement Learning (RL)
-Reinforcement learning seeks to find an optimal policy that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$.
+Reinforcement learning seeks to find an optimal policy, $\pi^{*}$.
+that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig.
+1-(a)).
 In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 19].
@@
 $\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
+(a) RL
+(b) IRL
+(c) GAIL + AIRL (terms in green)
+(d) RILe
+Figure 1: Overview of the related works.
+(a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data.
+IRL has two stages: 1.
+training a policy with frozen reward function, and 2.
+updating the reward function by comparing the converged policy with data.
+These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function.
+GAIL trains both policy and the discriminator at the same time.
+AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green).
+(D) RILe: similar to IRL, learning a reward function from data.
+RILe learns the reward function at the same time with the policy, using discriminator as a guide for learning the reward.
 3.3 Inverse Reinforcement Learning (IRL)
-Given sample trajectories $\tau_{E}$ of an optimal expert policy $\pi_{E}$, inverse reinforcement learning $\text{IRL}(\tau_{E})$ seeks to recover a reward function $R_{\theta}^{*}(s,a)$ that maximally rewards behavior that was sampled from the expert policy $\pi_{E}$.
-In other words, IRL aims to find a reward function $R_{\theta}^{*}(s,a)$ that satisfies
-$E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ (2)
-When IRL recovers the optimal reward function, $R_{\theta}^{*}(s,a)$ would result in a policy that exhibits expert behavior when optimized with reinforcement learning: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{argmax}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})].$
-Since only the trajectories created by the expert policy are observed, expectations are estimated from the samples in $\tau_{E}$.
-In other words, IRL seeks to find the reward function in which the expert policy performs better than any other policy, where the expectation of the expert is estimated through samples.
-With entropy regularization $H(\pi)$, maximum casual entropy inverse reinforcement learning [15] can be defined as
+Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig.
+1-(b)).
+Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]$.
+Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$.
+Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [14] is defined as
 $\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(E_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(E_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)
-3.4 Adversarial Imitation Learning (AIL)
-In contrast to inverse reinforcement learning, imitation learning aims to directly approximate the expert policy from given expert trajectory samples. It can be formulated as
+3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
+Imitation Learning (IL) aims to directly approximate the expert policy from given expert trajectory samples $\tau_{E}$. It can be formulated as
 $\text{IL}(\tau_{E})=\operatorname*{arg\,min}_{\pi}E_{(s,a)\sim\tau_{E}}[L(\pi(\cdot|s),a)],$ (3)
 where $L$ is a loss function, that captures the difference between policy and expert data.
-GAIL [3] extends imitation learning to an adversarial setting by quantifying the difference between policies of the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$.
-The discriminator’s function is to differentiate between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$.
-Its optimization problem is defined as
-$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\notin\tau_{E}}[\text{log}(1-D_{\phi}(s,a))].$ (5)
-Using the discriminator, the goal of GAIL is to find the optimal policy that minimizes this difference metric while maximizing an entropy constraint by training the discriminator and the policy at the same time.
-The optimization problem can be formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$, represented by
+GAIL [3] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)).
+The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$.
+The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint.
+The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
 $\min_{\pi}\max_{\phi}E_{\pi}[\text{log}\ D_{\phi}(s,a)]+E_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (4)
 In other words, the reward function that is maximized by the policy is defined as a similarity function, expressed as $R(s,a)=-\text{log}\ (D_{\phi}(s,a))$.
+AIRL [6] extends this adversarial approach to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig.
+1-(c)).
+The key insight of AIRL is to structure the discriminator as:
+$D_{\phi,\psi}(s,a,s^{\prime})=\frac{\exp(f_{\phi}(s,a,s^{\prime}))}{\exp(f_{\phi}(s,a,s^{\prime}))+\pi(a|s)},$ (5)
+where $f_{\phi}(s,a,s^{\prime})=r_{\psi}(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$.
+Here, $r_{\psi}(s,a)$ represents the learned reward function that is decoupled from the environment dynamics, $\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$.
+The AIRL optimization problem is formulated equivalently to GAIL (see Eqn.
+4).
+The reward function $r_{\psi}(s,a)$ is learned through minimizing the cross-entropy loss inherent in this adversarial setup.
+Therefore, the reward function remains tightly coupled with the discriminator’s learning process.
 4 RILe: Reinforced Imitation Learning
 We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process.
-Our framework consists of three key components: a student agent, a discriminator, and a teacher agent (Figure 1).
-Intuitively, the student agent learns the action policy and the teacher agent learns the reward function.
+Our framework consists of three key components: a student agent, a trainer agent, and a discriminator (Fig.
+2).
+Intuitively, the student agent learns the action policy and the trainer agent learns the reward function.
 Both agents are trained simultaneously via RL with the help of an adversarial discriminator.
-The discriminator enables training the policy and reward function in one learning process and permits replicating expert behavior through similarity rather than supervised state action mapping.
-However, uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning.
-This setup enables the teacher agent to learn a reward function separately by continuously adapting to changes in the student’s policy and guiding the student based on the discriminator’s feedback.
-As a result, the teacher can understand different regions of the state-action space through student’s interactions, effectively guiding it beyond the potentially noisy and imperfect expert state-action pairs.
+Building on AIL, we utilize the discriminator to enable joint training of policy and reward function.
+Unlike traditional IRL, the discriminator provides immediate feedback to the trainer, eliminating the need to wait for the student agent’s convergence.
+The trainer agent, trained via RL, is rewarded by the discriminator, and thus motivated to maximize the cumulative reward from the discriminator.
+The trainer learns to craft a reward function that promotes expert-like behavior in the student, even in complex scenarios.
+Crucially, the trainer’s RL-based approach optimizes for long-term performance rather than immediate feedback.
+This allows the trainer to encourage exploration and gradual improvement in three main ways.
+First, the trainer can associate its reward-giving actions with future improvements in the student’s behavior, even if those improvements take many steps to manifest.
+Second, the trainer can provide small positive rewards for actions that move the student in the right direction, even if they don’t immediately result in expert-like behavior.
+Finally, the trainer can adapt its reward function based on the student’s current capabilities, providing a learning path that gradually guides the student towards expert behavior.
+This approach enables RILe to overcome limitations of previous methods, particularly in complex tasks requiring extensive exploration, by promoting the discovery of expert-like strategies even when the student’s initial policy significantly diverges from expert behavior.
 In the following, we define the components of RILe and explain how they can efficiently learn behavior from imperfect data.
 Figure 2: Framework overview.
-The framework consists of three key components: a student agent, a teacher agent, and a discriminator.
-(1) The student agent learns a policy $\pi_{S}$ by interacting with an environment.
-(2) Based on the state action pair of the student agent $s^{T}=(s^{S},a^{S})$, the teacher agent learns a reward function as a policy $\pi_{T}$ by (3) choosing an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$.
-(4) The teacher agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator.
-(5) The discriminator compares student state-action pair with expert demonstrations $s^{D}$ and (6) the rewards the teacher based on their similarity $r^{T}=D(s^{T})$.
+The framework consists of three key components: a student agent, a trainer agent, and a discriminator.
+The student agent learns a policy $\pi_{S}$ by interacting with an environment, and the trainer agent learns a reward function as a policy $\pi_{T}$.
+(1) The student receives the environment state $s^{S}$.
+(2) The student takes an action $a^{S}$, forwards it to the environment which is updated based on $a^{S}$.
+(3) The student forwards its state and action to the trainer, whose state is $s^{T}=(s^{S},a^{S})$.
+(4) Trainer, $\pi_{T}$, evaluates the state action pair of the student agent $s^{T}=(s^{S},a^{S})$ and chooses an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$.
+(5) The trainer agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator.
+(6) Discriminator compares student state-action pair with expert demonstrations ($s^{D}$).
+(7) Discriminator gives reward to the trainer, based on the similarity between student- and expert-behavior.
 Student Agent
@@
 For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$.
-However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the teacher agent $\pi_{T}$.
-Therefore, the reward function is represented by the teacher policy.
-Thus, the student agent is guided by the actions of the teacher agent, i.e., the action of the teacher is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$.
+However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the trainer agent $\pi_{T}$.
+Therefore, the reward function is represented by the trainer policy.
+Thus, the student agent is guided by the actions of the trainer agent, i.e., the action of the trainer is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$.
 The optimization problem of the student agent is then defined as
@@
 The according lemma and proof are presented in the Appendix A.
-Teacher Agent
-The teacher agent aims to guide the student to approximate expert behavior by operating as its reward mechanism.
-Since the teacher agent takes the role of a reward function for the student, a new MDP is defined for the teacher agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1.
+Trainer Agent
+The trainer agent aims to guide the student to approximate expert behavior by operating as its reward mechanism.
+Since the trainer agent takes the role of a reward function for the student, a new MDP is defined for the trainer agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1.
 $A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value.
@@
 $K_{T}$ is the initial state distribution.
-The teacher agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$.
-Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the teacher comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$.
+The trainer agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$.
+Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the trainer comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$.
 It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1.
-To help the teacher understand how its actions impact the reward it receives, we define the reward function such that it multiplies the discriminator’s output by a sigmoid function dependent on the teacher’s actions.
-Therefore, the teacher agent’s reward function is defined as $R^{T}=\text{log}(D_{\phi}(s^{T}))\sigma(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\sigma(a^{T})$ is the sigmoid function applied to the teacher’s own actions.
-Considering only the discriminator’s output in the teacher agent’s reward function would result in a reward independent of the teacher’s actions.
-The optimization problem of the teacher can be defined as
-$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\text{log}(D_{\phi}(s,a))\sigma(a^{T})].$ (9)
+To help the trainer understand how its actions impact the reward it receives, we define the reward function such that it multiplies the scaled discriminator’s output by trainer’s actions.
+Therefore, the trainer agent’s reward function is defined as $R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\upsilon(x)=2x-1$ is the scaling function.
+By incorporating $a^{T}$ into the reward function, the trainer learns to adjust its policy based on the effectiveness of its previous actions.
+The optimization problem of the trainer can be defined as
+$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
 RILe
 RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$.
-In RILe, the student policy $\pi_{S}$ and the teacher policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method.
-The training algorithm is given in Appendix C.
+In RILe, the student policy $\pi_{S}$ and the trainer policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method.
+The training algorithm is given in Appendix D.
 Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as
 $\pi_{S}^{*}=\operatorname*{arg\,max}_{\pi_{S}}E_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)\right].$ (9)
-At the same time, the teacher agent aims to recover the optimal policy $\pi_{T}^{*}$ as
-$\pi_{T}^{*}=\operatorname*{argmax}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\text{log}(D_{\phi}(s^{T}_{t}))\sigma(a^{T})\right].$ (11)
-To prove that the student agent can learn expert-like behavior, we need to show that the teacher agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
-Lemma 1: Given the discriminator $D_{\phi}$, the teacher agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
+At the same time, the trainer agent aims to recover the optimal policy $\pi_{T}^{*}$ as
+$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}\right].$ (10)
+To prove that the student agent can learn expert-like behavior, we need to show that the trainer agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
+Lemma 1: Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
 Proof for Lemma 1 The assumptions are presented in Appendix A.
 The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$.
-The role of the teacher is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
+The role of the trainer is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
 We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student.
-The teacher’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
-The Value and Q functions of the teacher, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards.
-The value function for the teacher’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
+The trainer’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
+The Value and Q functions of the trainer, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards.
+The value function for the trainer’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
 $V^{\theta_{T}}(s_{t})=\mathbb{E}_{\pi^{\theta_{T}}}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}\right],$ (11)
-where $r_{t+k}=D_{\phi}(s_{t+k})$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
+where $r_{t+k}=\upsilon(D_{\phi}(s_{t+k}))a_{t+k}$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
 $Q^{\theta_{T}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{T}}}\left[V^{\theta_{T}}(s_{t+1})\mid s_{t},a_{t}\right].$ (12)
-The teacher’s policy optimization is done by maximizing the following clipped surrogate objective function:
+The trainer’s policy optimization is done by maximizing the following clipped surrogate objective function:
 $\begin{split}L^{CLIP}(\theta_{T})=&\mathbb{E}_{(s_{t},a_{t})\sim\pi^{\theta_{T}}}\Big{[}\min\Big{(}\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})}A^{\theta_{T_{old}}}(s_{t},a_{t}),\\ &\qquad\qquad\qquad\text{clip}(\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon)A^{\theta_{T_{old}}}(s_{t},a_{t})\Big{)}\Big{]},\end{split}$ (13)
-where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the teacher’s old policy parameters $\theta_{T_{old}}$.
+where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the trainer’s old policy parameters $\theta_{T_{old}}$.
 By expressing the advantage using the reward from the discriminator, we explicitly tie the policy gradient updates to the discriminator’s output, emphasizing the shaping of $\pi^{\theta_{S}}$ to match $p_{\text{expert}}(s,a)$:
 $A^{\theta_{T_{old}}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}(s_{t+1})\right]-V^{\theta_{T_{old}}}(s_{t}).$ (14)
-$A^{\theta_{T_{old}}}((s^{S}_{t},a^{S}_{t}),a_{t})=D_{\phi}((s^{S}_{t},a^{S}_{t}))+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}((s^{S}_{t+1},a^{S}_{t+1}))\right]-V^{\theta_{T_{old}}}(s^{S}_{t},a^{S}_{t}).$ (16)
 During each policy update, the objective in Equation 13 is maximized, driving parameter updates to favor actions that elicit higher rewards from the discriminator – effectively the actions that better align with expert behavior:
 $\theta_{T}\leftarrow\theta_{T}+\alpha_{\pi}\nabla_{\theta_{T}}L^{CLIP}(\theta_{T}).$ (15)
-The update rule in Equation 17, driven by cumulative rewards from the discriminator, incrementally adapts the teacher’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$.
-The teacher’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
+The update rule in Equation 15, driven by cumulative rewards from the discriminator, incrementally adapts the trainer’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$.
+The trainer’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
 5 Experiments
-We evaluate the performance of RILe by addressing three key questions.
-First, can RILe recover a feasible policy when trained to imitate from noisy expert data?
-To answer this question, we compare RILe’s performance with common imitation learning baselines in a grid world setting where policies are learned from noisy expert data.
-Second, can RILe use expert-data explicitly to imitate expert behavior?
-We use MuJoCo [25, 26] to evaluate RILe’s performance when expert data is leaked to the agents.
-Third, is RILe efficient and scalable to high-dimensional continuous control tasks?
-We answer this by using LocoMujoco [27] as the benchmark for evaluating RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks.
-Baselines We compared RILe with four baseline methods: Behavioral cloning (BC [7, 28]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [4]) and inverse reinforcement learning (IQ-Learn [20]).
-5.1 Noisy Expert Data
-(a) Expert traj.
-(b) RILe traj.
-(c) GAIL traj.
-(d) AIRL traj.
-(e) IQLearn traj.
-(f) RILe val.
-(g) GAIL val.
-(h) AIRL val.
-(i) IQLearn val.
-Figure 2: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death.
-(c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish.
-(b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal.
-Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively.
-The optimal path, derived from the reward of the teacher or discriminator, is highlighted with green lines.
-To demonstrate the advantage of using RL to learn the reward function in RILe, as opposed to deriving the reward directly from the discriminator in AIL and AIRL, we designed a 5x5 MiniGrid experiment.
-The grid consists of 4 lava tiles that immediately kill the agent if it steps in it, representing terminal conditions.
-The goal condition of the environment is reaching the green tile.
-The expert demonstrations are imperfect, depicting an expert that passes through a lava tile without being killed and still reaches the green goal tile.
-Using this data, we trained the adversarial approaches with a perfect discriminator, which provides a reward of $0.99$ if the visited state-action pair stems from the expert and $0.01$ otherwise.
-These values were chosen over 1 and 0 because both AIRL and GAIL use the logarithm of the discriminator output to calculate rewards.
-Results are presented in Fig.
-2.
-The value graphs (Fig.
-2e-g) are attained by computing the value of each grid cell $c_{i}$ as $\sum_{a\in A}D(c_{i},a)$ for AIRL and GAIL, and $\sum_{a\in A}\pi_{T}(c_{i},a)$ for RILe.
+We evaluate the performance of RILe by addressing three key questions:
+  1. 1.
+    How does RILe’s adaptive reward function evolve compared to baseline methods and how does this evolution enhance the learning process?
+  2. 2.
+    Is RILe efficient and scalable to high-dimensional continuous control tasks?
+  3. 3.
+    Can RILe use expert-data explicitly to imitate expert behavior?
+To answer the first question, we compare RILe’s performance with AI(R)L baselines in a maze setting, where we demonstrate how the trainer agent modifies the reward function to guide the student during training.
+For the second question, we evaluate RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks, using LocoMujoco [20] as the environment.
+To answer the last question, we use a simulated humanoid character from MuJoCo [21, 22] to evaluate RILe’s performance when expert data is explicitly used by the agents.
+Additional experimental results are provided in the Appendix, where we evaluate the robustness of the learned reward function in changing environments and analyze the noise resilience of our method.
+Baselines We compare RILe with four baseline methods: Behavioral cloning (BC [7, 23]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [6]) and inverse reinforcement learning (IQ-Learn [24]).
+5.1 Evolving Reward Function
+To evaluate the impact of RILe’s trainer agent on the learning process compared to existing methods, we designed a maze experiment.
+The task is to navigate from a fixed start on the left side to a goal in the top-right corner, avoiding obstacles.
+Using a single expert demonstration, we trained RILe, GAIL, and AIRL.
 Fig.
-2a shows the expert trajectory.
-GAIL (Fig.
-2c), AIRL (Fig.
-2d) and IQLearn (Fig.
-2e) fail to reach the goal, as their agents either become stuck or are directed into lava.
-In contrast, RILe (Fig.
-2d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data.
-The difference in the value graphs between RILe and the baselines intuitively explains this outcome.
-In AIL and AIRL (Fig.
-2f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly.
-Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory.
-In contrast, RILe’s teacher agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting.
-By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig.
-2f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal.
-Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration.
-Since the student agent is guided by the teacher to also match the expert trajectory, it remains close to this path after passing the lava tiles.
-5.2 Expert Data Leakage in Training
-We investigate whether RILe can use expert data explicitly to imitate expert behavior better.
-Thus, we analyze how the performance and convergence time will be effected when we leak expert data to the replay buffers of both student and teacher agent.
-For this experiment, we employed MuJoCo’s Humanoid environment [26, 25] and used a single trajectory from the expert data of [20].
-The expert data is leaked to agents’ replay buffers with following percentages: $25\%,50\%,75\%,100\%$.
-$25\%$ signifies that $25\%$ of the replay buffer’s data is sourced from the expert trajectory, while the remaining $75\%$ originates from the agents’ interactions with the environment.
-$100\%$ indicates that the agents’ environment interactions are entirely disregarded, with only expert data being used for training.
-As shown in Figure 3, leaking expert data to buffers slightly worsened performance but significantly reduced the convergence time.
-However, when environment interactions were entirely replaced with expert data, the policy’s performance declined substantially.
-For comparison, results from IQLearn and Behavioral Cloning are also presented, as both methods rely exclusively on expert data for training.
-Neither method, however, achieved the same level of performance as RILe.
-Normalized Scores
-Normalized Steps
-Figure 3: Scores (left) representing the total rewards achieved, and convergence steps (right) representing the number of steps to convergence, are presented against the percentage of leaked expert data.
-For RILe, the x-axis shows the percentage of leaked expert data in the student agent’s buffer, while the y-axis shows it in the teacher agent’s buffer.
-Scores/Steps are normalized by no-leakage training results of RILe.
-Behavioral Cloning fails to achieve a comparable score to RILe.
-While IQLearn converges faster than the no-leakage RILe, it still falls short in terms of overall performance.
+3 shows how the student’s reward function evolves during training with all three methods.
+For RILe, we plot the reward function learned by the trainer.
+For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator.
+The columns represent reward plots at 25%, 50%, 75%, and 100% of training completion for each algorithm, respectively.
+We present the student agent’s trajectory from the previous epoch on reward plots to demonstrate how reward functions adapt to the student agent’s progress.
+(a) RILe
+(b) GAIL
+(c) AIRL
+Figure 3: Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment.
+Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right).
+The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black.
+Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards.
+Purple squares represent obstacles.
+RILe demonstrates a more adaptive reward function that evolves with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
+The discriminator-based reward functions of GAIL and AIRL remain relatively static, largely independent of the student agent’s current performance.
+This static behavior is because the discriminators in GAIL and AIRL focus on distinguishing expert data and progressively improve throughout training.
+In contrast, the trainer in RILe adapts itself based on the current learning progress of the student’s policy.
+As the first column shows, when the student’s trajectory does not resemble the expert, RILe’s trainer encourage exploration towards the expert path.
+This indicates that the trainer can provide informative rewards despite potentially low discriminator feedback.
+Then, in the second column, as the student progresses, RILe’s reward function adapts dynamically.
+When the student learns to reach the bottom-right, the trainer shifts high rewards to the top-left, guiding the agent to explore that area.
+Third column shows as the student approaches the goal, the trainer increases rewards around it while maintaining rewards in specific areas (e.g., the left part) to prevent the agent from getting stuck.
+Finally, as the fourth column shows, the trainer focuses high rewards on the top-right area near the goal while maintaining a path of rewards that guides the student along the expert trajectory.
+RILe’s evolving reward function demonstrates its ability to provide meaningful guidance even when the discriminator easily identifies non-expert policies.
+By adapting to the student’s current capabilities, RILe creates a dynamic learning curriculum that encourages exploration and gradual improvement towards expert-like behavior.
+Our results demonstrate that the trainer effectively associates rewards with future behavior improvements, even if delayed.
+It provides incremental rewards for actions moving the student in the right direction, and adapts the reward function to match the student’s abilities, gradually guiding them toward expert behavior.
 5.2 Motion-Capture Data Imitation for Robotic Continuous Control
-In LocoMujoco, the objective is to learn policies that can accurately imitate the collected MoCap data of various robotic tasks.
-LocoMujoco is challenging due its high dimensionality.
-The MoCap data only consists of states of the robot, not its actions, which prevents the use of Behavioral Cloning and IQLearn in this analysis, since they both requires to have full expert data with states and actions.
-It should be noted that LocoMujoco environments do not have terminal or goal conditions.
-Results are presented in Table 1 for both test seeds and validation seeds. RILe outperforms baselines especially for new initial conditions introduced by varying seeds.
+We evaluate RILe’s performance on the LocoMujoco benchmark, which presents a challenging task of imitating motion-capture (MoCap) data for various robotic control tasks.
+This benchmark is particularly demanding due to its high dimensionality and the absence of action data in the MoCap recordings which precludes the use of methods like Behavioral Cloning and IQLearn that require complete state-action pairs.
 Table 1: Test and validation results on seven LocoMujoco tasks.
@@
 Talos Carry 220.1 186.3 134.2 503.3 338.5 74.1 1000
-UnitreeH1 Walk 898.3 950.2 568.1 980.7 965.1 716.2 1000
+UnitreeH1 Walk 966.2 950.2 568.1 980.7 965.1 716.2 1000
 UnitreeH1 Carry 788.3 634.6 130.5 850.6 637.4 140.9 1000
 Humanoid Walk 831.3 181.4 80.1 970.3 216.2 78.2 1000
+Table 1 presents the results for seven LocoMujoco tasks, comparing RILe with GAIL and AIRL across both test and validation seeds.
+The expert performance, set at 1000 for all tasks, serves as the upper benchmark.
+RILe demonstrates superior performance in all scenarios, particularly excelling in generalization to new initial conditions as evidenced by the test seed results.
+5.3 Impact of Expert Data on Trainer-Student Dynamics
+Figure 4: Impact of explicit usage of expert data on RILe’s performance.
+Red markers and yellow markers show normalized scores and normalized steps, respectively.
+Square markers show IQ-Learn and BC results for comparisons.
+The percentage of expert data in the buffers are presented around the figure.
+More expert data in buffers speeds convergence but can reduce final performance, especially at 100%, showing how the explicit usage of excess expert data hinders trainer-student adaptation in RILe
+We study how explicitly incorporating expert data into RILe’s training affects the trainer’s ability to adapt to the student’s needs, analyzing performance and convergence time when expert data is added to the agents’ replay buffers.
+For this experiment, we used MuJoCo’s Humanoid environment [22, 21] and a single expert trajectory from [24].
+We varied the proportion of expert data in the replay buffers from 0% to 100%; for example, 25% means a quarter of the buffer is expert data and 75% is from agent interactions.
+As shown in Fig.
+4, introducing expert data led to faster convergence but decreased performance.
+Notably, when environmental interactions were completely replaced by the expert data (100% case), the student’s performance declined significantly.
+Excessive expert data appears to hinder the trainer’s ability to adapt to the student, disrupting RILe’s dynamic learning process.
+For comparison, we include results from IQ-Learn and Behavioral Cloning (BC), which rely explicitly on expert data.
+Despite their exclusive use of expert data, neither method matched RILe’s performance, even when RILe used substantial amounts of expert data.
+Our findings underscore the importance of balancing environmental interactions and expert demonstrations in buffers in RILe.
+While some expert data accelerates learning, too much impedes the trainer’s adaptability, underscoring the unique strength of RILe’s interactive learning approach.
 6 Discussion
-We have demonstrated through our experiments that RILe outperforms baseline models across various settings.
-While RILe surpasses the state-of-the-art adversarial IL/IRL methods, it also has its limitations.
-The main challenge is training the policy with a changing reward function, which leads to unstable policy updates for the student agent.
-To address this issue, we formulate the teacher’s reward by correlating its actions with the output of the discriminator, instead of only relying on the discriminator.
-However, instability remains a challenge and causes slower learning of the action policy in RILe compared to AIL.
-Future work should establish bounds for the reward agent updates to enhance the stability of the student agent’s learning process.
-Moreover, balancing the learning rates of the discriminator and the policies is challenging, particularly because the discriminator tends to overfit the data quickly.
-This overfitting makes it difficult for the teacher agent to find a reward for the student that can challenge the discriminator.
-While adjusting the discriminator’s learning rate on a case-by-case basis helps, future research should focus on a fundamental solution, such as replacing the discriminator with a more robust similarity component.
-Uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning.
-Therefore, the reward function is learned, not directly derived from the discriminator.
-This distinction allows RILe to provide meaningful guidance even when the discriminator overfits to noisy expert data.
-In the LocoMujoco benchmark, where perfect expert data is available, we argue that RILe’s ability to account for the agent’s exploration efforts during training, in addition to its behavioral similarity with the expert, results in a superior generalization capability compared to AIL and AIRL.
+Our experiments have demonstrated that RILe consistently outperforms baseline models across various settings, surpassing state-of-the-art adversarial IL/IRL methods.
+RILe’s key strength lies in its adaptive learning approach, where the trainer agent dynamically adjusts the reward function based on the student’s current capabilities.
+This enables more efficient learning, particularly in complex environments.
+However, RILe faces challenges, primarily in maintaining policy stability with a changing reward function.
+Our solution of freezing the trainer after the student achieves half the expert score is effective but potentially limiting.
+Additionally, the discriminator’s tendency to overfit quickly presents another significant challenge.
+Future work should focus on establishing bounds for reward agent updates to enhance stability without compromising adaptability, developing a stronger theoretical understanding of RILe’s convergence properties, and exploring discriminator-less approaches.
+Despite these challenges, RILe has demonstrated advantages in adaptability, robustness, and generalization, thereby making it a promising direction for future research in imitation learning and opening up new possibilities for dynamic and responsive learning processes.
 Appendix A Justification of RILe
@@
   • The discriminator loss curve is complex and the discriminator function, $D_{\phi}(s,a)$, is sufficiently expressive since it is parameterized by a neural network with adequate capacity.
-  • For the teacher’s and student’s policy functions $(\pi^{\theta_{T}})$ and $(\pi^{\theta_{S}})$, and the Q-functions $(Q^{\theta_{S}})$, each is Lipschitz continuous with respect to its parameters with constants $(L_{\theta_{T}}),(L_{\theta_{S}}),and(L_{Q})$, respectively.
+  • For the trainer’s and student’s policy functions $(\pi^{\theta_{T}})$ and $(\pi^{\theta_{S}})$, and the Q-functions $(Q^{\theta_{S}})$, each is Lipschitz continuous with respect to its parameters with constants $(L_{\theta_{T}}),(L_{\theta_{S}}),and(L_{Q})$, respectively.
 This means for all $(s,a)$ and for any pair of parameter settings $(\theta,\theta^{\prime}):[|\pi^{\theta}(s,a)-\pi^{\theta^{\prime}}(s,a)|\leq L_{\theta}|\theta-\theta^{\prime}|,][|Q^{\theta}(s,a)-Q^{\theta^{\prime}}(s,a)|\leq L_{Q}|\theta-\theta^{\prime}|.]$
@@
 The discriminator’s ability to distinguish between student and expert pairs improves as $\mathcal{L}_{D}(\phi)$ is minimized, implying that $\lim_{n_{batch}\rightarrow\infty}D{\phi^{*}}(s,a)=P(\pi_{E}|s,a)$, where $n_{batch}$ is the number of batches.
-Appendix B Compute Resources
+Appendix B Additional Experiments
+B.1 Noisy Expert Data
+(a) Expert traj.
+(b) RILe traj.
+(c) GAIL traj.
+(d) AIRL traj.
+(e) IQLearn traj.
+(f) RILe val.
+(g) GAIL val.
+(h) AIRL val.
+(i) IQLearn val.
+Figure 5: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death.
+(c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish.
+(b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal.
+Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively.
+The optimal path, derived from the reward of the trainer or discriminator, is highlighted with green lines.
+To demonstrate the advantage of using RL to learn the reward function in RILe, as opposed to deriving the reward directly from the discriminator in AIL and AIRL, we designed a 5x5 MiniGrid experiment.
+The grid consists of 4 lava tiles that immediately kill the agent if it steps in it, representing terminal conditions.
+The goal condition of the environment is reaching the green tile.
+The expert demonstrations are imperfect, depicting an expert that passes through a lava tile without being killed and still reaches the green goal tile.
+Using this data, we trained the adversarial approaches with a perfect discriminator, which provides a reward of $0.99$ if the visited state-action pair stems from the expert and $0.01$ otherwise.
+These values were chosen over 1 and 0 because both AIRL and GAIL use the logarithm of the discriminator output to calculate rewards.
+Results are presented in Fig.
+5.
+The value graphs (Fig.
+5e-g) are attained by computing the value of each grid cell $c_{i}$ as $\sum_{a\in A}D(c_{i},a)$ for AIRL and GAIL, and $\sum_{a\in A}\pi_{T}(c_{i},a)$ for RILe.
+Fig.
+5a shows the expert trajectory.
+GAIL (Fig. 5c), AIRL (Fig. 5d) and IQLearn (Fig. 5e) fail to reach the goal, as their agents either become stuck or are directed into lava.
+In contrast, RILe (Fig.
+5d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data.
+The difference in the value graphs between RILe and the baselines intuitively explains this outcome.
+In AIL and AIRL (Fig.
+5f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly.
+Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory.
+In contrast, RILe’s trainer agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting.
+By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig.
+5f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal.
+Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration.
+Since the student agent is guided by the trainer to also match the expert trajectory, it remains close to this path after passing the lava tiles.
+B.2 Robustness to Noise in the Expert Data
+To evaluate the robustness of RILe and baseline methods to noise in the expert data, we conducted experiments in the MuJoCo Humanoid-v2 environment.
+Artificial noise sampled from a zero-mean Gaussian distribution with varying standard deviations ($\Sigma$) was added to a single expert trajectory, affecting either the actions or the states.
+The baselines used for comparison were GAIL [3], AIRL [6], RIL-Co [25], IC-GAIL [26], and IQ-Learn [24].
+As shown in Table 2, RILe consistently outperforms the baselines across different noise levels, demonstrating superior robustness even when a high amount of noise is present in the expert data ($\Sigma=0.5$).
+These results indicate that RILe is less sensitive to imperfections in the expert demonstrations compared to existing methods.
+Table 2: Test results in MuJoCo Humanoid-v2 environment, where artificial noise sampled from a zero-mean Gaussian distribution is added to a single expert trajectory.
+Results are aggregated over 20 different-seed environments.
+IQ-Learn* is trained using the official code and hyperparameters of the IQ-Learn algorithm.
+Noise-Free Action Noise State Noise
+$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$ $\Sigma=0.2$ $\Sigma=0.5$
+RILe 5681 5280 5154 5350 5205
+GAIL 5430 5275 902 5147 917
+AIRL 5276 4869 4589 4898 4780
+RIL-Co 576 491 493 505 501
+IC-GAIL 610 601 568 590 591
+IQ-Learn* 312 192 153 243 277
+B.3 Robustness of the Learned Reward Function
+We evaluated the robustness of the reward functions learned by RILe and AIRL [6] through an experiment similar to that conducted by [27].
+Initially, both methods were trained to learn reward functions in a noise-free MuJoCo Humanoid-v2 environment.
+After training, these reward functions were frozen.
+Subsequently, new student agents were trained using these fixed reward functions in environments where Gaussian noise was added to the agents’ actions, with varying noise levels.
+Table 3 presents the results of this evaluation.
+The reward function learned by RILe demonstrates superior robustness to noise, maintaining high performance even under increased noise levels.
+In contrast, the performance of agents using the reward function learned by AIRL decreases more significantly as noise increases.
+These findings indicate that the reward function learned by RILe is more resilient to environmental noise, contributing to better agent performance in noisy conditions.
+Table 3: We test the robustness of learned reward functions.
+After training reward functions in a noise-free setting, reward functions are frozen, and used to train a new agent in a noisy environment, where Gaussian noise is added to agent’s actions in every step.
+No Noise Mild Noise High Noise
+$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$
+RILe 5748 5201 5196
+AIRL 5334 5005 4967
+B.4 Reward Curves
+We compare the reward curves of RILe, GAIL [3], AIRL [6], IQ-Learn [24], and AdapMen [28] in the MuJoCo Humanoid-v2 experiment.
+Since the task involves learning from expert trajectories, we combined AdapMen with an adversarial discriminator to enable training without an expert policy.
+As shown in the reward curves, despite RILe having multiple components, it is the most efficient method.
+This efficiency is achieved through the dynamic guidance of the trainer during training, which adapts the reward function to meet the student’s needs.
+Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen.
+AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
+Appendix C Compute Resources
 For the training of RILe and baselines, following computational sources are employed:
   • AMD EPYC 7742 64-Core Processor
-  • 1 x Nvidia RTX6000 GPU
+  • 1 x Nvidia A100 GPU
   • 32GB Memory
-Appendix C Algorithm
+Appendix D Algorithm
 Algorithm 1 RILe Training Process
-1:Initialize student policy $\pi_{S}$ and teacher policy $\pi_{T}$ with random weights, and the discriminator $D$ with random weights.
+1:Initialize student policy $\pi_{S}$ and trainer policy $\pi_{T}$ with random weights, and the discriminator $D$ with random weights.
 2:Initialize an empty replay buffer $B$
@@
 6:     for each transition $(s,a)$ in $\tau_{S}$ do
-7:         Calculate student reward $R^{S}$ using teacher policy:
+7:         Calculate student reward $R^{S}$ using trainer policy:
 $R^{S}\leftarrow\pi_{T}$ (19)
@@
 12:     for each transition $(s,a)$ in $\tau_{S}$ do
-13:         Calculate teacher reward $R^{T}$ using discriminator:
-$R^{T}\leftarrow\log(D(s,a))$ (23)
+13:         Calculate trainer reward $R^{T}$ using discriminator:
+$R^{T}\leftarrow\upsilon(D(s,a))a^{T}$ (21)
 14:         Update $\pi_{T}$ using policy gradient with reward $R^{T}$

### v3

@@
 Abstract
-Reinforcement Learning has achieved significant success in generating complex behavior but often requires extensive reward function engineering.
-Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator.
-However, these methods struggle in complex tasks where randomly sampling expert-like behaviors is challenging.
-This limitation stems from their reliance on policy-agnostic discriminators, which provide insufficient guidance for agent improvement, especially as task complexity increases and expert behavior becomes more distinct.
-We introduce RILe (Reinforced Imitation Learning environment), a novel trainer-student system that learns a dynamic reward function based on the student’s performance and alignment with expert demonstrations.
-In RILe, the student learns an action policy while the trainer, using reinforcement learning, continuously updates itself via the discriminator’s feedback to optimize the alignment between the student and the expert.
-The trainer optimizes for long-term cumulative rewards from the discriminator, enabling it to provide nuanced feedback that accounts for the complexity of the task and the student’s current capabilities.
-This approach allows for greater exploration of agent actions by providing graduated feedback rather than binary expert/non-expert classifications.
-By reducing dependence on policy-agnostic discriminators, RILe enables better performance in complex settings where traditional methods falter, outperforming existing methods by 2x in complex simulated robot-locomotion tasks.
+Acquiring complex behaviors is essential for artificially intelligent agents, yet learning these behaviors in high-dimensional settings poses a significant challenge due to the vast search space.
+Traditional reinforcement learning (RL) requires extensive manual effort for reward function engineering.
+Inverse reinforcement learning (IRL) uncovers reward functions from expert demonstrations but relies on an iterative process that is often computationally expensive.
+Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons offer insufficient feedback for effective learning.
+We introduce RILe (Reinforced Imitation Learning), a framework that combines the strengths of imitation learning and inverse reinforcement learning to learn a dense reward function efficiently and achieve strong performance in high-dimensional tasks.
+RILe employs a novel trainer–student framework: the trainer learns an adaptive reward function, and the student uses this reward signal to imitate expert behaviors.
+By dynamically adjusting its guidance as the student evolves, the trainer provides nuanced feedback across different phases of learning.
+Our framework produces high-performing policies in high-dimensional tasks where direct imitation fails to replicate complex behaviors.
+We validate RILe in challenging robotic locomotion tasks, demonstrating that it significantly outperforms existing methods and achieves near-expert performance across multiple settings.
 1 Introduction
-Reinforcement Learning (RL) has emerged as a powerful framework for teaching agents to perform complex tasks.
-In recent years, deep reinforcement learning has demonstrated remarkable success in replicating sophisticated behaviors, including playing Atari games, chess, and Go [1, 2].
-However, these achievements often come at a cost: the tedious and challenging process of designing reward functions, as predicting the policy outcome from a manually crafted reward function remains notoriously difficult.
-To overcome the reward engineering problem, Imitation Learning (IL) leverages expert demonstrations to learn a policy.
-Since vast amounts of expert data are required to accurately learn expert behaviors, Adversarial Imitation Learning (AIL) approaches, such as GAIL [3], have been proposed as data-efficient alternatives.
-AIL employs a discriminator to measure similarity between learned behavior and expert behavior, rewarding the agent accordingly.
-While computationally efficient, AIL methods suffer from a critical limitation: the policy-agnostic nature of their discriminators.
-The discriminator lacks any inherent incentive to guide the agent towards expert-like behavior, in contrast to engineered reward functions in RL.
-Consequently, AIL methods face challenges in complex tasks requiring extensive exploration to find optimal actions.
-For instance, in digital locomotion tasks, AIL methods often struggle to consistently replicate expert performance [4].
-Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering.
-Unlike IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that motivates the agent to acquire expert behaviors.
-The reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior.
-This iterative process renders IRL computationally expensive [5].
-Adversarial Inverse Reinforcement Learning (AIRL) [6] attempts to address this inefficiency by introducing a discriminator that enables simultaneous learning of the policy and reward function.
-However, in AIRL, the reward function is tightly coupled to the discriminator, potentially limiting its ability to capture complex task structures or long-term dependencies and inheriting the limitations of a policy-agnostic discriminators.
-This highlights the need for a method that can learn a more flexible reward function without the computational overhead of traditional IRL methods.
-To overcome these challenges and effectively learn behaviors in complex settings, we propose Reinforced Imitation Learning (RILe) (Fig.
+Learning complex behaviors is critical for advancing artificially intelligent agents in fields such as robotics and strategic games.
+Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design, which is both time-consuming and error-prone.
+There are two ways to address the reward engineering problem.
+First, Inverse Reinforcement Learning (IRL) [1, 2] offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering.
+IRL proceeds iteratively: it first trains a policy (the learning agent’s decision-making mechanism) using the current reward function, observes how well the agent’s behavior aligns with the expert’s, and then refines the reward function to better guide the policy toward expert-like behaviors.
+Repeating this process eventually yields a reward function capable of providing nuanced feedback at different stages of learning.
+However, this iterative procedure is computationally expensive [3], especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
+Second, Imitation learning (IL) bypasses explicit reward design by directly comparing learned behaviors to expert demonstrations via a comparison mechanism.
+Traditional IL approaches such as Behavioral Cloning (BC) [4] match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks.
+To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL [5], introduce a discriminator that operates as a comparison mechanism and judges how expert-like the learned behaviors are.
+However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements.
+Instead, they rely on a comparison metric, often a distance measure or binary classification, that merely checks whether the agent’s behavior is (or is not) similar to the expert.
+Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize.
+Consequently, both traditional IL and AIL struggle in high-dimensional environments [6, 7], where the agent needs more granular and adaptive feedback than these mechanisms provide.
+Adversarial Inverse Reinforcement Learning (AIRL) [8] attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator.
+However, AIRL tightly couples the reward function to the discriminator’s output, causing it to inherit AIL’s limitations in high-dimensional settings where more fine-grained guidance is needed.
+Real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog.
+The teacher also refines how they teach as the student progresses.
+Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student’s learning progress more effectively.
+By contrast, existing approaches lack this continuous cooperative synergy.
+Adversarial Imitation Learning (AIL) does update a discriminator alongside the policy, but the discriminator’s sole role is to distinguish expert-like behavior from non-expert behavior.
+Consequently, the student attempts to fool this judge into classifying its behavior as expert-like, resulting in a competitive process rather than a cooperative trainer that dynamically shapes rewards based on suboptimal behaviors.
+Meanwhile, IRL methods only refine the reward after the policy converges, missing the opportunity for continuous co-evolution throughout training.
+To address these issues, inspiring by these insights, we propose Reinforced Imitation Learning (RILe).
+RIle combines the adaptive reward benefits of IRL with the computational efficiency of AIL (Fig.
 1-(d)).
-RILe aims to combine the ability to learn a reward function that actively guides the agent to imitate expert behavior with the computational efficiency of adversarial frameworks.
-At the core of RILe is a novel trainer-student system designed to address the shortcomings of existing methods:
-  • A student agent that learns to replicate the expert’s policy via RL in the environment
-  • A trainer agent that learns a reward function via RL and guides the student agent during training
-By integrating the trainer-student dynamic, RILe decouples reward learning from policy learning and the discriminator, allowing each component to specialize and thereby overcome the limitations of policy-agnostic discriminators.
-While RILe utilizes a discriminator similar to those in adversarial frameworks, its role is fundamentally redefined.
-In RILe, the discriminator’s primary function is to provide feedback to the trainer agent by distinguishing expert data from student roll-outs.
-This feedback serves as the reward signal for the trainer, not directly influencing the student agent.
-The trainer leverages the discriminator’s feedback to learn a reward function that effectively guides the student agent.
-This approach enables more nuanced reward shaping, particularly beneficial in tasks requiring complex decision-making and extensive exploration.
+RILe is a novel trainer-student system that establishes a positive-sum relationship between the trainer and the student.
+Specifically, RILe is composed of:
+  • Student Agent: Learns a policy to imitate expert demonstrations using reinforcement learning.
+  • Trainer Agent: Simultaneously learns a reward function using reinforcement learning, leveraging an adversarial discriminator for continuous feedback on student performance.
+RILe’s trainer continuously updates the reward function in tandem with the student’s policy updates, where IRL refines its reward function only after training a policy to convergence.
+Specifically, the trainer queries a discriminator to measure how expert-like the student’s behavior is, then optimizes the reward function based on that feedback, without waiting for the policy to converge.
+RILe preserves nuanced reward shaping, while avoiding IRL’s heavy computational loop.
+As a result, RILe is particularly effective in high-dimensional settings, where agents need fine-grained guidance at every stage of learning.
 Our contributions are two-fold:
   1. 1.
-    Decoupled Reward-function Learning: We introduce a novel approach where the trainer agent learns the reward function independently from both the student agent and the discriminator.
-Unlike existing methods that derive rewards directly from discriminator outputs, our trainer agent uses reinforcement learning to optimize the reward function based on the feedback from the discriminator.
-By focusing on long-term reward maximization, RL enables the trainer to distill inconsistent feedback from the discriminator into meaningful rewards, leading to better student performance.
+    Efficient Reward-Function Learning via RL: We introduce a reinforcement-learning-based approach for training a reward function simultaneously with the policy.
+This avoids IRL’s repeated policy re-training and the purely discriminator-based rewards of AIL/AIRL.
+Reinforcement learning enables the trainer agent to explore multiple reward regimes, yielding a reward function that considers long-horizon effects.
   2. 2.
-    Dynamic Reward Customization: Our trainer agent dynamically adjusts rewards based on the student agent’s progress, facilitating a better learning experience and enabling accurate imitation of expert behavior in complex settings.
-This adaptive approach allows for more gradual learning, particularly in tasks where the optimal behavior may change depending on the agent’s current capabilities.
-We evaluate RILe against state-of-the-art methods in AIL, and AIRL, specifically GAIL [3], and AIRL [6].
-Our experiments span three distinct scenarios: (1) Tailoring a reward function dynamically in a discrete maze task, (2) Investigating the impact of expert data on the trainer-student dynamics in a humanoid locomotion task, and (3) Imitating motion-capture data in continuous control tasks.
-The results demonstrate RILe’s superior performance, especially in complex tasks, and its ability to learn an effective dynamic reward function where baseline methods fail.
+    Dynamic Reward Customization: RILe offers context-sensitive guidance at every stage of training, because the trainer agent updates the reward function as the student evolves.
+This dynamic shaping is especially valuable in high-dimensional tasks, where the learning agent requires different forms of encouragement during intermediate-stages than later-stages of the training.
+Consequently, RILe enables accurate imitation of the expert performance in high-dimensional tasks.
+We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL, specifically GAIL [5] AIRL [8], GAIfO [9], BCO [10], IQ-Learn [7] and DRAIL [11].
+Our experiments spans four scenarios: (1) Empirically investigating how RILe’s reward-learning strategy differs from AIL and AIRL in a maze task, (2) Quantifying the adaptability of the learned reward function in a humanoid locomotion task, (3) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems, and (4) Analyzing the impact of explicit expert-data usage within RILe’s training process.
+Our results show RILe’s superior performance, particularly in high-dimensional environments, and highlight RILe’s ability to learn a dynamic reward function that effectively guides the student through multiple stages of training.
 2 Related Work
-We review literature on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), which form the conceptual foundation of RILe.
+We review research on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), the conceptual foundations of RILe.
 Imitation Learning
-Early work introduced Behavioral Cloning (BC) [7], which learns a policy congruent with expert demonstrations through supervised learning.
-DAgger [8] introduces data aggregation.
-GAIL [3] introduces adversarial methods, where a discriminator aims to discriminate expert demonstrations, while a generator tries to fool the discriminator.
-DQfD [9] proposes two-stage approach with pre-training, and ValueDice [10] uses a distribution-matching objective between policy and expert.
-Despite progress, IL faces challenges in efficacy and generalization [5, 11].
-RILe addresses these by introducing an adaptive teacher agent to guide the student beyond expert demonstrations.
+Early work in IL introduced Behavioral Cloning (BC) [4], which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations.
+DAgger [12] refines BC by aggregating data over multiple iterations to mitigate compounding errors.
+GAIL [5] employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior.
+BCO [10] extends BC, and GAIfO [9] extends GAIL, both to handle state-only observation scenarios.
+DQfD [13] introduces a two-stage approach with pre-training, while ValueDice [14] aligns policy and expert distributions via a distribution-matching objective.
+More recently, DRAIL [11] leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
+Despite these advances, IL methods face challenges in high-dimensional environments [6, 7], where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance.
+RILe addresses these limitations through an adaptive trainer–student framework, where a learned reward function provides more nuanced guidance than standard IL comparison mechanisms.
 Inverse Reinforcement Learning
-IRL, introduced by [12], learns the expert’s intrinsic reward function.
-Key developments include Apprenticeship Learning [13], Maximum Entropy IRL [14], and adversarial approaches like AIRL [6].
-Recent work explores handling unstructured data [15] and cross-embodiment scenarios [16].
-Despite advancements, IRL faces challenges in computational efficiency and scalability [17].
-RILe addresses these by jointly learning policy and reward function in a single process.
+Inverse Reinforcement Learning (IRL), introduced by [1], aims to uncover the expert’s intrinsic reward function from demonstrations.
+Major developments include Apprenticeship Learning [15], Maximum Entropy IRL [2], and adversarial variants like AIRL [8].
+IQ-Learn [7] reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability.
+More recent work focuses on unstructured data [16] and cross-embodiment transfer [17].
+Nonetheless, IRL methods struggle with computational inefficiency and limited scalability [18], particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly.
+RILe mitigates these challenges by jointly learning the policy and reward function in a single process, avoiding IRL’s iterative retraining loop and facilitating more efficient reward shaping for complex environments.
 3 Background
@@
 A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$.
-$S$ is the state space that consists of all possible environment states $s$, and $A$ is action space that contains all possible environment actions $a$.
+$S$ is the state space consisting of all possible environment states $s$, and $A$ is action space containing all possible environment actions $a$.
 $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function.
-$T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$.
+$T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$.
 $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor.
@@
 In this work, we consider $\gamma$-discounted infinite horizon settings.
-Following [3], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $E_{\pi}[R(s,a)]\triangleq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$.
+Following [5], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$.
 The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$.
 We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [19].
-Our work considers an imitation learning problem from expert trajectories, each consisting of states $s$ and actions $a$.
+Our work considers an imitation learning problem from expert trajectories, consisting of states $s$ and actions $a$.
 The set of expert trajectories $\tau_{E}$ are sampled from an expert policy $\pi_{E}\in\Pi$, where $\Pi$ is the set of all possible policies.
@@
 1-(a)).
-In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 19].
-The RL problem with a parameterized reward function and entropy regularization is defined as
-$\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
+In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ [5, 20].
+The RL problem with a parameterized reward function and entropy regularization is defined as:
+$\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
 (a) RL
@@
 (D) RILe: similar to IRL, learning a reward function from data.
-RILe learns the reward function at the same time with the policy, using discriminator as a guide for learning the reward.
+RILe learns the reward function at the same time with the policy, using a discriminator as a guide for learning the reward.
 3.3 Inverse Reinforcement Learning (IRL)
@@
 1-(b)).
-Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]$.
+Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $\mathbb{E}_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})+H(\pi)]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}$.
 Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$.
 Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [2] is defined as
-$\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(E_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(E_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)
+$\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(\mathbb{E}_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)
 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
-Imitation Learning (IL) aims to directly approximate the expert policy from given expert trajectory samples $\tau_{E}$. It can be formulated as
-$\text{IL}(\tau_{E})=\operatorname*{arg\,min}_{\pi}E_{(s,a)\sim\tau_{E}}[L(\pi(\cdot|s),a)],$ (3)
-where $L$ is a loss function, that captures the difference between policy and expert data.
-GAIL [3] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)).
-The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$.
+[…]
+GAIL [5] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig.
+1-(c)).
+The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert ones $(s,a)\notin\tau_{E}$.
 The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint.
 The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
-$\min_{\pi}\max_{\phi}E_{\pi}[\text{log}\ D_{\phi}(s,a)]+E_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (4)
+$\min_{\pi}\max_{\phi}\mathbb{E}_{\pi}[\text{log}\ D_{\phi}(s,a)]+\mathbb{E}_{\tau_{E}}[\text{log}\ (1-D_{\phi}(s,a))]-\lambda H(\pi).$ (3)
 In other words, the reward function that is maximized by the policy is defined as a similarity function, expressed as $R(s,a)=-\text{log}\ (D_{\phi}(s,a))$.
-AIRL [6] extends this adversarial approach to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig.
-1-(c)).
-The key insight of AIRL is to structure the discriminator as:
+AIRL [8] extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as:
 $D_{\phi,\psi}(s,a,s^{\prime})=\frac{\exp(f_{\phi}(s,a,s^{\prime}))}{\exp(f_{\phi}(s,a,s^{\prime}))+\pi(a|s)},$ (4)
@@
 The AIRL optimization problem is formulated equivalently to GAIL (see Eqn.
-4).
+3).
 The reward function $r_{\psi}(s,a)$ is learned through minimizing the cross-entropy loss inherent in this adversarial setup.
@@
 4 RILe: Reinforced Imitation Learning
-We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process.
-Our framework consists of three key components: a student agent, a trainer agent, and a discriminator (Fig.
-2).
-Intuitively, the student agent learns the action policy and the trainer agent learns the reward function.
-Both agents are trained simultaneously via RL with the help of an adversarial discriminator.
-Building on AIL, we utilize the discriminator to enable joint training of policy and reward function.
-Unlike traditional IRL, the discriminator provides immediate feedback to the trainer, eliminating the need to wait for the student agent’s convergence.
-The trainer agent, trained via RL, is rewarded by the discriminator, and thus motivated to maximize the cumulative reward from the discriminator.
-The trainer learns to craft a reward function that promotes expert-like behavior in the student, even in complex scenarios.
-Crucially, the trainer’s RL-based approach optimizes for long-term performance rather than immediate feedback.
-This allows the trainer to encourage exploration and gradual improvement in three main ways.
-First, the trainer can associate its reward-giving actions with future improvements in the student’s behavior, even if those improvements take many steps to manifest.
-Second, the trainer can provide small positive rewards for actions that move the student in the right direction, even if they don’t immediately result in expert-like behavior.
-Finally, the trainer can adapt its reward function based on the student’s current capabilities, providing a learning path that gradually guides the student towards expert behavior.
-This approach enables RILe to overcome limitations of previous methods, particularly in complex tasks requiring extensive exploration, by promoting the discovery of expert-like strategies even when the student’s initial policy significantly diverges from expert behavior.
-In the following, we define the components of RILe and explain how they can efficiently learn behavior from imperfect data.
-Figure 2: Framework overview.
+We propose Reinforced Imitation Learning (RILe) to jointly learn a reward function and a policy that emulates expert-like behavior within a single learning process.
+RILe introduces a novel trainer–student dynamic, as illustrated in Figure 2.
+In RILe, the student agent learns an action policy by interacting with the environment, while the trainer agent learns a reward function that effectively guides the student toward expert-like behavior.
+Both agents are trained simultaneously via reinforcement learning, with an assistance from an adversarial discriminator.
+Specifically, the trainer queries the discriminator, which judges how expert-like the student’s behavior is, and then optimizes the reward function based on that feedback on-the-fly.
+Unlike traditional AIL, where the discriminator effectively is employed as the reward function for the student, RILe introduces a trainer agent to provide fine-grained feedback to the student, while avoiding IRL’s iterative computational expense.
+The trainer agent plays the key role in RILe.
+Trained via RL, the trainer explores different reward designs and learns to provide gradually tailored feedback to the student by maximizing the cumulative rewards it receives from the discriminator.
+This approach equips RILe with two key advantages over existing IRL/AIRL/AIL frameworks: (1) On-the-fly reward function learning via RL: The reward function is learned continuously with RL, enabling the trainer to explore different reward options and account for long-horizon effects of its signals, (2) Context-sensitive guidance: The trainer adjusts its reward outputs in response to the student’s current policy, thereby encouraging the student to explore suboptimal actions that ultimately guide it closer to expert behavior.
+By providing tailored feedback at different stages of training, RILe addresses the limitations of prior methods, particularly in high-dimensional tasks.
+In the remainder of this section, we define the components of RILe and explain how they jointly learn from expert demonstrations.
+Figure 2: Reinforced Imitation Learning (RILe).
 The framework consists of three key components: a student agent, a trainer agent, and a discriminator.
@@
 Student Agent
-The student agent aims to learn a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP.
+The student agent learns a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP.
 For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$.
-However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the trainer agent $\pi_{T}$.
+However, instead of using a handcrafted reward function, the student’s reward comes from the trainer agent’s policy, $\pi_{T}$.
 Therefore, the reward function is represented by the trainer policy.
@@
 The optimization problem of the student agent is then defined as
-$\min_{\pi_{S}}-E_{(s^{S},a^{S})\sim\pi_{S}}[\pi_{T}\left((s^{S},a^{S})\right)].$ (6)
+$\min_{\pi_{S}}-\mathbb{E}_{(s^{S},a^{S})\sim\pi_{S}}[\pi_{T}\left((s^{S},a^{S})\right)].$ (5)
 Discriminator
-The discriminator differentiates between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and state-action pairs stemming from the policy of the student $(s,a)\sim\pi_{S}$.
+The discriminator differentiates between expert-generated state-action pairs, $(s,a)\sim\tau_{E}$, and pairs from the student, $(s,a)\sim\pi_{S}$.
 In RILe, the discriminator is defined as a feed-forward deep neural network, parameterized by $\phi$.
-Hence, the optimization problem is
-$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\sim\pi_{S}}[\text{log}(1-D_{\phi}(s,a))].$ (7)
-To provide effective guidance, the discriminator needs to accurately distinguish whether a given state-action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$.
-The feasibility of this discrimination has been demonstrated by GAIL [3].
-The according lemma and proof are presented in the Appendix A.
+Its objective is:
+$\max_{\phi}\mathbb{E}_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+\mathbb{E}_{(s,a)\sim\pi_{S}}[\text{log}(1-D_{\phi}(s,a))].$ (6)
+To provide effective guidance, the discriminator must accurately identify whether a given state–action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$.
+GAIL [5] established the feasibility of such a discriminator (see Appendix B for details).
 Trainer Agent
-The trainer agent aims to guide the student to approximate expert behavior by operating as its reward mechanism.
-Since the trainer agent takes the role of a reward function for the student, a new MDP is defined for the trainer agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1.
-$A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value.
-$R_{T}(s,a)$ is the reward function where $s\in S_{T}$ and $a\in A_{T}$.
-$T_{T}=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as the state distribution upon taking action $a\in A_{T}$ in state $s\in S_{T}$.
-$K_{T}$ is the initial state distribution.
-The trainer agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$.
-Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the trainer comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$.
-It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1.
-To help the trainer understand how its actions impact the reward it receives, we define the reward function such that it multiplies the scaled discriminator’s output by trainer’s actions.
-Therefore, the trainer agent’s reward function is defined as $R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\upsilon(x)=2x-1$ is the scaling function.
-By incorporating $a^{T}$ into the reward function, the trainer learns to adjust its policy based on the effectiveness of its previous actions.
-The optimization problem of the trainer can be defined as
-$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
+The trainer agent guides the student toward expert behavior by serving as its reward mechanism.
+Since the trainer does not directly observe the student’s policy $\pi_{S}$, we model the trainer’s environment as a Partially Observable MDP (POMDP): $\text{POMDP}_{T}=(S_{T},A_{T},\Omega_{T},T_{T},O_{T},R_{T},\gamma)$.
+The state space $S_{T}=S\times A\times\pi_{S}$ includes all possible state-action pairs from the standard MDP and the student’s policy $\pi_{S}$, which is hidden from the trainer, introducing partial observability.
+The trainer’s action space, $A_{T}$, consists of scalar values.
+Formally, $A_{T}$ is defined as a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$.
+The observation space $\Omega_{T}=S\times A$ consists of the observable state-action pairs of the student.
+The transition dynamics $T_{T}$ and the observation function $O_{T}$ are defined formally in Appendix A.
+The reward function $R_{T}(s^{T},a^{T})$ evaluates the effectiveness of the trainer’s action in guiding the student, where $s^{T}=(s^{S},a^{S})$ is the observation of the trainer.
+$\gamma$ is the discount factor.
+Within this POMDP, the trainer learns a policy $\pi_{T}$ that produces helpful reward signals for $\pi_{S}$.
+The trainer observes only the student’s state–action pair, $s^{T}=(s^{S},a^{S})\in S\times A$, not $\pi_{S}$ itself.
+It then outputs a scalar action $a^{T}\in[-1,1]$, which is provided to the student as the reward, $r^{S}$.
+If the trainer’s reward depends only on the discriminator’s output, the trainer receives the same reward regardless of whether it rewards or penalizes the student, yielding no immediate feedback on its choices.
+For instance, if the student behaves like the expert and the discriminator outputs $\approx$ 1, the trainer should ideally reward the student (action, $a^{T}$, $\approx$ 1).
+But if the trainer’s action is not factored into its own reward, it gains no immediate signal whether rewarding or punishing the student was effective, since it receives the same reward in either case.
+This ambiguity forces extensive trial and error.
+To address this, we define the trainer reward as:
+$R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$ (7)
+where $\upsilon(x)=2x-1$ scales the discriminator’s output, making it symmetric around zero.
+Including $a^{T}$ in the trainer’s reward ensures the trainer effectively learns from its own actions.
+Formally, the trainer’s objective is:
+$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
 RILe
-RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$.
-In RILe, the student policy $\pi_{S}$ and the trainer policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method.
-The training algorithm is given in Appendix D.
-Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as
-$\pi_{S}^{*}=\operatorname*{arg\,max}_{\pi_{S}}E_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)\right].$ (9)
-At the same time, the trainer agent aims to recover the optimal policy $\pi_{T}^{*}$ as
-$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}E_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}\right].$ (10)
-To prove that the student agent can learn expert-like behavior, we need to show that the trainer agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
-Lemma 1: Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
-Proof for Lemma 1 The assumptions are presented in Appendix A.
-The expert’s state-action distribution is denoted by $p_{\text{expert}}(s,a)$.
-The role of the trainer is to provide a reward signal to the student that encourages the approximation of $p_{\text{expert}}(s,a)$ as closely as possible.
-We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student.
-The trainer’s policy $\pi^{\theta_{T}}$, parameterized by $\theta_{T}$, aims to maximize the likelihood under the discriminator’s assessment, thus encouraging the student agent to generate state-action pairs drawn from a distribution resembling $p_{\text{expert}}(s,a)$.
-The Value and Q functions of the trainer, conditioned on the rewards provided by the discriminator, are defined in terms of expected cumulative discriminator rewards.
-The value function for the trainer’s policy parameters $\theta_{T}$ at state $s_{t}$, simplified from $s^{T}_{t}$ for better readability, is given by:
-$V^{\theta_{T}}(s_{t})=\mathbb{E}_{\pi^{\theta_{T}}}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}\mid s_{t}\right],$ (11)
-where $r_{t+k}=\upsilon(D_{\phi}(s_{t+k}))a_{t+k}$. Similarly, the Q-function for taking action $a_{t}$ in state $s_{t}$ and then following policy $\pi^{\theta_{T}}$ can be written as:
-$Q^{\theta_{T}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{T}}}\left[V^{\theta_{T}}(s_{t+1})\mid s_{t},a_{t}\right].$ (12)
-The trainer’s policy optimization is done by maximizing the following clipped surrogate objective function:
-$\begin{split}L^{CLIP}(\theta_{T})=&\mathbb{E}_{(s_{t},a_{t})\sim\pi^{\theta_{T}}}\Big{[}\min\Big{(}\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})}A^{\theta_{T_{old}}}(s_{t},a_{t}),\\ &\qquad\qquad\qquad\text{clip}(\frac{\pi^{\theta_{T}}(a_{t}|s_{t})}{\pi^{\theta_{T_{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon)A^{\theta_{T_{old}}}(s_{t},a_{t})\Big{)}\Big{]},\end{split}$ (13)
-where $A^{\theta_{T_{old}}}(s_{t},a_{t})$ is the advantage function computed as $Q^{\theta_{T_{old}}}(s_{t},a_{t})-V^{\theta_{T_{old}}}(s_{t})$, defined with respect to the trainer’s old policy parameters $\theta_{T_{old}}$.
-By expressing the advantage using the reward from the discriminator, we explicitly tie the policy gradient updates to the discriminator’s output, emphasizing the shaping of $\pi^{\theta_{S}}$ to match $p_{\text{expert}}(s,a)$:
-$A^{\theta_{T_{old}}}(s_{t},a_{t})=D_{\phi}(s_{t})+\gamma\mathbb{E}_{\pi^{\theta_{S}}}\left[V^{\theta_{T_{old}}}(s_{t+1})\right]-V^{\theta_{T_{old}}}(s_{t}).$ (14)
-During each policy update, the objective in Equation 13 is maximized, driving parameter updates to favor actions that elicit higher rewards from the discriminator – effectively the actions that better align with expert behavior:
-$\theta_{T}\leftarrow\theta_{T}+\alpha_{\pi}\nabla_{\theta_{T}}L^{CLIP}(\theta_{T}).$ (15)
-The update rule in Equation 15, driven by cumulative rewards from the discriminator, incrementally adapts the trainer’s policy to reinforce student behaviors that are indistinguishable from those of the expert according to $D_{\phi}$.
-The trainer’s policy is guided to facilitate the student’s approximation of $p_{\text{expert}}(s,a)$, thereby fulfilling its role in the imitation learning process.
+RILe brings together these three components, student, trainer, and discriminator, to discover a student policy that imitates expert behaviors in $\tau_{E}$.
+Both $\pi_{S}$ and $\pi_{T}$ can be trained via any single-agent RL method.
+The overall training algorithm is detailed in Appendix J.
+The student agent aims to recover the optimal policy $\pi_{S}^{*}$:
+$\pi_{S}^{*}=\operatorname*{arg\,max}_{\pi_{S}}\mathbb{E}_{(s^{S},a^{S})\sim\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\pi_{T}\left((s^{S}_{t},a^{S}_{t})\right)]\right].$ (9)
+Simultaneously, the trainer aims to recover $\pi_{T}^{*}$:
+$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}_{t}]\right].$ (10)
+By optimizing these objectives together, RILe efficiently learns both a reward function and a policy in high-dimensional settings where traditional AIL or IRL methods often struggle.
+Details on specific training strategies are provided in Appendix C.
 5 Experiments
-We evaluate the performance of RILe by addressing three key questions:
+We evaluate the performance of RILe by addressing four key questions:
   1. 1.
-    How does RILe’s adaptive reward function evolve compared to baseline methods and how does this evolution enhance the learning process?
+    How does RILe’s reward-learning strategy differs from AIL and AIRL?
   2. 2.
-    Is RILe efficient and scalable to high-dimensional continuous control tasks?
+    How adaptive is RILe’s learned reward function?
   3. 3.
-    Can RILe use expert-data explicitly to imitate expert behavior?
-To answer the first question, we compare RILe’s performance with AI(R)L baselines in a maze setting, where we demonstrate how the trainer agent modifies the reward function to guide the student during training.
-For the second question, we evaluate RILe’s effectiveness in imitating motion-capture data within continuous robotic control tasks, using LocoMujoco [20] as the environment.
-To answer the last question, we use a simulated humanoid character from MuJoCo [21, 22] to evaluate RILe’s performance when expert data is explicitly used by the agents.
-Additional experimental results are provided in the Appendix, where we evaluate the robustness of the learned reward function in changing environments and analyze the noise resilience of our method.
-Baselines We compare RILe with four baseline methods: Behavioral cloning (BC [7, 23]), adversarial imitation learning (GAIL [3]), adversarial inverse reinforcement learning (AIRL [6]) and inverse reinforcement learning (IQ-Learn [24]).
+    How does RILe perform in both low- and high-dimensional continuous-control problem compared to AIL/AIRL/IRL?
+  4. 4.
+    How explicitly using expert-data within RILe’s training process affects the context-sensitive reward learning?
+Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC [4, 21], BCO [10]), adversarial imitation learning (GAIL [5], GAIfO [9] and DRAIL [11]), adversarial inverse reinforcement learning (AIRL [8]), and inverse reinforcement learning (IQ-Learn [7]).
+DRAIL [11] introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
+Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix H.
 5.1 Evolving Reward Function
-To evaluate the impact of RILe’s trainer agent on the learning process compared to existing methods, we designed a maze experiment.
-The task is to navigate from a fixed start on the left side to a goal in the top-right corner, avoiding obstacles.
-Using a single expert demonstration, we trained RILe, GAIL, and AIRL.
-Fig.
-3 shows how the student’s reward function evolves during training with all three methods.
-For RILe, we plot the reward function learned by the trainer.
-For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator.
-The columns represent reward plots at 25%, 50%, 75%, and 100% of training completion for each algorithm, respectively.
-We present the student agent’s trajectory from the previous epoch on reward plots to demonstrate how reward functions adapt to the student agent’s progress.
+To answer the first question about RILe’s reward-learning strategy, we compare RILe’s performance with AI(R)L baselines in a maze setting.
+In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
 (a) RILe
@@
 (c) AIRL
-Figure 3: Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment.
+Figure 3: Reward Function Comparison.
+Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment.
 Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right).
@@
 Purple squares represent obstacles.
-RILe demonstrates a more adaptive reward function that evolves with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
-The discriminator-based reward functions of GAIL and AIRL remain relatively static, largely independent of the student agent’s current performance.
-This static behavior is because the discriminators in GAIL and AIRL focus on distinguishing expert data and progressively improve throughout training.
-In contrast, the trainer in RILe adapts itself based on the current learning progress of the student’s policy.
-As the first column shows, when the student’s trajectory does not resemble the expert, RILe’s trainer encourage exploration towards the expert path.
-This indicates that the trainer can provide informative rewards despite potentially low discriminator feedback.
-Then, in the second column, as the student progresses, RILe’s reward function adapts dynamically.
-When the student learns to reach the bottom-right, the trainer shifts high rewards to the top-left, guiding the agent to explore that area.
-Third column shows as the student approaches the goal, the trainer increases rewards around it while maintaining rewards in specific areas (e.g., the left part) to prevent the agent from getting stuck.
-Finally, as the fourth column shows, the trainer focuses high rewards on the top-right area near the goal while maintaining a path of rewards that guides the student along the expert trajectory.
-RILe’s evolving reward function demonstrates its ability to provide meaningful guidance even when the discriminator easily identifies non-expert policies.
-By adapting to the student’s current capabilities, RILe creates a dynamic learning curriculum that encourages exploration and gradual improvement towards expert-like behavior.
-Our results demonstrate that the trainer effectively associates rewards with future behavior improvements, even if delayed.
-It provides incremental rewards for actions moving the student in the right direction, and adapts the reward function to match the student’s abilities, gradually guiding them toward expert behavior.
-5.2 Motion-Capture Data Imitation for Robotic Continuous Control
-We evaluate RILe’s performance on the LocoMujoco benchmark, which presents a challenging task of imitating motion-capture (MoCap) data for various robotic control tasks.
-This benchmark is particularly demanding due to its high dimensionality and the absence of action data in the MoCap recordings which precludes the use of methods like Behavioral Cloning and IQLearn that require complete state-action pairs.
-Table 1: Test and validation results on seven LocoMujoco tasks.
-Test Seeds Validation Seeds
-RILe GAIL AIRL RILe GAIL AIRL Expert
-Atlas Walk 870.6 792.7 300.5 895.4 918.6 356 1000
-Atlas Carry 850.8 669.3 256.4 889.7 974.2 271.9 1000
-Talos Walk 842.5 442.3 102.1 884.7 675.5 103.4 1000
-Talos Carry 220.1 186.3 134.2 503.3 338.5 74.1 1000
-UnitreeH1 Walk 966.2 950.2 568.1 980.7 965.1 716.2 1000
-UnitreeH1 Carry 788.3 634.6 130.5 850.6 637.4 140.9 1000
-Humanoid Walk 831.3 181.4 80.1 970.3 216.2 78.2 1000
-Table 1 presents the results for seven LocoMujoco tasks, comparing RILe with GAIL and AIRL across both test and validation seeds.
-The expert performance, set at 1000 for all tasks, serves as the upper benchmark.
-RILe demonstrates superior performance in all scenarios, particularly excelling in generalization to new initial conditions as evidenced by the test seed results.
-5.3 Impact of Expert Data on Trainer-Student Dynamics
-Figure 4: Impact of explicit usage of expert data on RILe’s performance.
-Red markers and yellow markers show normalized scores and normalized steps, respectively.
-Square markers show IQ-Learn and BC results for comparisons.
-The percentage of expert data in the buffers are presented around the figure.
-More expert data in buffers speeds convergence but can reduce final performance, especially at 100%, showing how the explicit usage of excess expert data hinders trainer-student adaptation in RILe
-We study how explicitly incorporating expert data into RILe’s training affects the trainer’s ability to adapt to the student’s needs, analyzing performance and convergence time when expert data is added to the agents’ replay buffers.
-For this experiment, we used MuJoCo’s Humanoid environment [22, 21] and a single expert trajectory from [24].
-We varied the proportion of expert data in the replay buffers from 0% to 100%; for example, 25% means a quarter of the buffer is expert data and 75% is from agent interactions.
-As shown in Fig.
-4, introducing expert data led to faster convergence but decreased performance.
-Notably, when environmental interactions were completely replaced by the expert data (100% case), the student’s performance declined significantly.
-Excessive expert data appears to hinder the trainer’s ability to adapt to the student, disrupting RILe’s dynamic learning process.
-For comparison, we include results from IQ-Learn and Behavioral Cloning (BC), which rely explicitly on expert data.
-Despite their exclusive use of expert data, neither method matched RILe’s performance, even when RILe used substantial amounts of expert data.
-Our findings underscore the importance of balancing environmental interactions and expert demonstrations in buffers in RILe.
-While some expert data accelerates learning, too much impedes the trainer’s adaptability, underscoring the unique strength of RILe’s interactive learning approach.
+RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
+Figure 3 shows how each method’s learned reward function evolves during training.
+For RILe, we plot the trainer’s learned reward function.
+For GAIL and AIRL, we visualize the discriminator outputs.
+The columns represent reward landscapes at 25%, 50%, 75%, and 100% of the total training process, and each subplot overlays the student’s trajectory from the previous epoch.
+RILe’s reward function dynamically adapts to the student’s current policy, providing guidance that encourage suboptimal actions which eventually lead the student closer to the expert trajectory.
+By contrast, GAIL and AIRL’s reward functions remain relatively static.
+Specifically, the first column in Figure 3 shows RILe’s trainer encouraging exploration toward the bottom-right of the maze, which is initially suboptimal but helpful in the long run.
+As the student learns to reach the lower part of the maze, RILe shifts high-reward regions toward the top-left (second column), again encouraging incremental progress.
+The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around other key areas to keep the agent from getting stuck.
+Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance.
+This dynamic reward adaptation gets important in higher-dimensional tasks as we show in Section 5.3.
+5.2 Reward Function Dynamics
+To address the second question on adaptability, we quantify how the reward function evolves and its relationship to the student’s performance.
+We compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
+We introduce three metrics (see Appendix D.2 for details): (1) RFDC (Reward Function Distribution Change): Wasserstein distance between reward distributions over consecutive training intervals, capturing overall shifts in reward space, (2) FS-RFDC (Fixed-State Reward Function Distribution Change): Mean absolute deviation of reward values at a fixed set of expert states over time, (3) CPR (Correlation between Performance and Reward): Pearson correlation between changes in the reward function and changes in the student’s performance.
+(a) RFDC
+(b) FS-RFDC
+(c) CPR
+Figure 4: Dynamics of Reward Functions.
+(a) Reward Function Distribution Change (RFDC): Wasserstein distance between reward function distributions.
+(b) Fixed-State Reward Function Distribution Change (FS-RFDC): Mean absolute deviation of reward values for a fixed set of expert states.
+(c) Correlation between Performance and Reward (CPR): Pearson correlation between changes in the reward function and changes in the student’s performance.
+5.2.1 Adaptability of the Learned Reward Function
+We assess how adaptive the reward function learned by RILe is compared to that of AIL.
+Fig.
+4(a) presents changes in reward distributions over 10,000 consecutive steps.
+RILe exhibits the highest adaptability in its reward function, aligning with our goal of having the reward function adapt based on the student’s learning stage.
+The advanced discriminator in DRAIL reduces the need for drastic reward function changes, yet RILe remains more adaptive than GAIL.
+Since the changing student policy indirectly affects the Fig.
+4(a), we also show changes in reward values for the fixed set of states in Fig.
+4(b).
+Again, RILe’s reward function is the most adaptive among all methods.
+Higher adaptability of RILe ensures that the reward signal remains aligned with incremental performance improvements, enabling the student to receive more timely and effective guidance throughout training.
+5.2.2 Correlation between the Learned Reward and the Student Performance
+We evaluate how changes in the reward function correlate with improvements in student performance.
+To this end, Fig.
+4(c) presents the Pearson correlation between student’s performance and reward updates.
+DRAIL-RILe achieves the highest positive correlation, indicating that it learns the most effective rewards for improving student performance.
+RILe ranks second, demonstrating that the trainer agent effectively helps the student achieve better scores even with the help of a naive-discriminator.
+In contrast, GAIL’s correlation starts positive but soon turns negative and remains so throughout training.
+We hypothesize that this occurs because the discriminator in GAIL tends to saturate as training progresses.
+While the discriminator’s reward signal effectively guides learning early on, its increasingly static nature at later stages fails to capture subtle performance improvements, leading to a negative correlation.
+5.3 Imitation in Continuous Control Tasks
+5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
+To answer the third question, we first test RILe in high-dimensional robotic locomotion tasks [22], where the agent must imitate motion-capture data for various robotic bodies.
+This benchmark is especially demanding due to its complexity and dimensionality.
+Table 1: Test results on seven LocoMujoco tasks.
+RILe GAIL AIRL IQ BCO GAIfO DRAIL GAIL DRAIL RILe Expert
+Walk Atlas 870.6 792.7 300.5 30.9 21.0 834.2 834.4 899.1 1000
+Talos 842.5 442.3 102.1 4.5 11.9 710.0 787.7 896.6 1000
+UnitreeH1 966.2 950.2 568.1 8.8 34.8 526.8 940.8 995.8 1000
+Humanoid 831.3 181.4 80.1 4.5 3.5 706.5 814.6 527.6 1000
+Carry Atlas 850.8 669.3 256.4 36.8 20.3 810.1 516.6 317.1 1000
+Talos 220.1 186.3 134.2 10.5 10.3 212.5 836.7 840.5 1000
+UnitreeH1 788.3 634.6 130.5 14.4 21.1 604.5 796.7 909.5 1000
+Table 1 presents results for seven LocoMujoco tasks across different test seeds (see Appendix D.3 for details).
+Overall, these results underscore that RILe outperforms the AIL/IL/IRL baselines, and benefits even more from the enhancements provided by the DRAIL variants, achieving performance levels close to the expert.
+The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role, suggesting avenues for future research.
+This overall performance aligns with our claim that an adaptive reward function is crucial for mastering complex, high-dimensional behaviors
+Table 2: Test results on four MuJoCo tasks.
+RILe GAIL AIRL IQ
+Humanoid 5928 5709 5623 327
+Walker2d 4435 4906 4823 270
+Hopper 3417 3361 3014 310
+HalfCheetah 5205 4173 3991 755
+Figure 5: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps, respectively. Expert data usage speeds the training of RILe but reduces final performance.
+5.3.2 Learning from Demonstrations
+We further evaluate RILe’s performance on four standard MuJoCo tasks [23, 24], Humanoid, Walker2d, Hopper, and HalfCheetah, with baseline comparisons widely reported in prior work.
+Table 2 summarizes test results.
+Overall, these results demonstrate that RILe consistently achieves competitive or superior performance compared to established AIL/AIRL/IRL methods.
+In particular, on high-dimensional tasks such as Humanoid, RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function in handling complex continuous control challenges.
+5.4 Explicit Usage of Expert Data
+To answer the fourth question, we use the Humanoid environment in MuJoCo [23, 24] with a single expert trajectory from [7].
+We vary the proportion of expert data in the replay buffers from 0% to 100%.
+For instance, 25% indicates a quarter of each buffer is expert data and 75% is agent data (see Appendix D.5 for details).
+Figure 5 illustrates how changing the proportion of expert data in both the trainer’s and the student’s replay buffers influences RILe’s training.
+Introducing more expert data accelerates RILe’s convergence (needs fewer training steps), but reduce the final performance.
+In the extreme case of 100% expert data, the student’s performance drops substantially.
+This indicates that too much expert data hamper the trainer’s ability to adapt to the student’s real-time needs, disrupting RILe’s context-sensitive reward customization.
+We also compare RILe’s results with IQ-Learn and BC, both of which rely heavily on expert data.
+Even when RILe is trained with substantial amounts of expert data, it still generally performs better, indicating that RILe’s adaptive reward-shaping provides a crucial edge over supervised methods.
 6 Discussion
-Our experiments have demonstrated that RILe consistently outperforms baseline models across various settings, surpassing state-of-the-art adversarial IL/IRL methods.
-RILe’s key strength lies in its adaptive learning approach, where the trainer agent dynamically adjusts the reward function based on the student’s current capabilities.
-This enables more efficient learning, particularly in complex environments.
-However, RILe faces challenges, primarily in maintaining policy stability with a changing reward function.
-Our solution of freezing the trainer after the student achieves half the expert score is effective but potentially limiting.
-Additionally, the discriminator’s tendency to overfit quickly presents another significant challenge.
-Future work should focus on establishing bounds for reward agent updates to enhance stability without compromising adaptability, developing a stronger theoretical understanding of RILe’s convergence properties, and exploring discriminator-less approaches.
-Despite these challenges, RILe has demonstrated advantages in adaptability, robustness, and generalization, thereby making it a promising direction for future research in imitation learning and opening up new possibilities for dynamic and responsive learning processes.
-Appendix A Justification of RILe
+As our experiments demonstrate, RILe consistently outperforms baseline models across various tasks, thanks to its adaptive learning approach, where the trainer agent continuously adjusts the reward based on the student’s current learning stage.
+Our maze experiments provide an interpretable example of how the trainer agent tailors its rewards.
+By encouraging actions that might seem suboptimal for immediate imitation but advantageous for long-term learning, RILe establishes a curriculum that ultimately boosts performance.
+This adaptive strategy helps RILe to achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
+Nonetheless, policy stability remains challenging when the reward function is constantly evolving.
+Freezing the trainer (see Appendix C) stabilizes learning but halts further adaptation, and the discriminator itself tends to overfit quickly.
+Future work could explore fully cooperative multi-agent RL techniques to allow ongoing adaptation, investigate ways to bound or regulate trainer updates, and consider discriminator-less formulations for reward learning.
+Despite these challenges, RILe shows that cooperatively learning the policy and the reward function can offer significant advantages over static or iteratively updated methods.
+By providing dynamic and tailored rewards, RILe effectively guides the student through complex tasks.
+We believe this opens up new possibilities for responsive and adaptive learning frameworks in imitation learning and beyond.
+Appendix A POMDP of the Trainer
+Partially Observable Markov Decision Process (POMDP) of the trainer is defined as $\text{POMDP}_{T}=(S_{T},A_{T},\Omega_{T},T_{T},O_{T},R_{T},\gamma)$.
+Here, $T_{T}=\{P(.\mid f^{T},a^{T})\}$ is the transition dynamics where $P(.\mid f^{T},a^{T})$ is the state distribution upon taking action $a\in A_{T}$ in state $f\in S_{T}$.
+The transition function incorporates the student’s policy $\pi_{S}$, which evolves in response to the rewards provided, reflecting the hidden dynamics due to the unobserved $\pi_{S}$.
+The observation function $O_{T}=\{P(s^{T}\mid f^{T},a^{T})\}$ defines the probability of observing $s^{T}\in\Omega_{T}$ given the state $(f^{T},a^{T})$.
+The trainer deterministically observes the student’s state-action pair, so $P(s^{T}=(s^{S},a^{S})\mid f^{T},a^{T})=1$, where $f^{T}=(s^{S},a^{S},\pi_{S})$.
+Appendix B Justification of RILe
 Assumptions:
@@
 This means for all $(s,a)$ and for any pair of parameter settings $(\theta,\theta^{\prime}):[|\pi^{\theta}(s,a)-\pi^{\theta^{\prime}}(s,a)|\leq L_{\theta}|\theta-\theta^{\prime}|,][|Q^{\theta}(s,a)-Q^{\theta^{\prime}}(s,a)|\leq L_{Q}|\theta-\theta^{\prime}|.]$
-A.1 Lemma 2:
+To prove that the student agent can learn expert-like behavior, we need to show that the trainer agent learns to give higher rewards to student experiences that match with the expert state-action pair distribution, as this would enable a student policy to eventually mimic expert behavior.
+B.1 Lemma 1:
+Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
+Proof for Lemma 1 The student agent, $\pi_{S}(a^{S}_{t}|s^{S}_{t})$, interacts with the environment and generates state-action pairs as $(s^{S}_{t},a^{S}_{t})$.
+The trainer agent observes these pairs and provides a reward $r^{S}_{t}=a^{T}_{t}=\pi_{T}(a^{T}_{t}|(s^{S}_{t},a^{S}_{t}))$ to the student, where $a^{T}_{t}\in[-1,1]$ is the trainer’s action.
+We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student.
+The trainer’s reward at timestep $t$ is:
+$r_{t}^{T}=\upsilon(D_{\phi}(s_{t}^{T}))a_{t}^{T}$ (11)
+where $s_{t}^{T}=(s^{S}_{t},a^{S}_{t})$ is the trainer’s observation, $D_{\phi}(s_{t}^{T})$ is the discrimantor output that estimates the likelihood that $s_{t}^{T}$ comes from the expert data, and $\upsilon(D)=2D-1$ is a scaling function that maps discriminator’s output to the range $[-1,1]$.
+The trainer maximizes the expected cumulative reward:
+$J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{T}\right]$ (12)
+where $\gamma\in[0,1)$ is the discount factor. In other words, trainer aims to find the policy that maximizes $J_{T}(\pi_{T})$: $\pi^{*T}=\operatorname*{arg\,max}_{\pi^{T}}J_{T}(\pi_{T})$.
+From the policy gradient theorem, the gradient of the trainer’s objective with respect to the policy parameters, $\theta_{T}$, is:
+$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})Q_{T}(s_{t}^{T},a_{t}^{T})\right]\$ (13)
+where $Q_{T}(s_{t}^{T},a_{t}^{T})$ is the action-value function of the trainer.
+The action-value function, $Q_{T}(s_{t}^{T},a_{t}^{T})$, and the value function, $V_{T}(s^{T}_{t})$ is defined by Bellman equation as:
+$Q_{T}(s_{t}^{T},a_{t}^{T})$ $=r_{t}^{T}+\gamma\mathbb{E}_{s^{T}_{t+1}}\left[V_{T}(s^{T}_{t+1})\right]$ (14)
+$V_{T}(s^{T}_{t+1})$ $=\mathbb{E}_{a_{t}^{T}\sim\pi_{T}}\left[Q_{T}(s_{t}^{T},a_{t}^{T}))\right]$ (15)
+The trainer aims to maximize $Q_{T}(s^{T}_{t},a^{T}_{t})$ to satisfy Equation 13.
+Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ has the same sign as $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
+$\alpha^{*T}_{t}=\begin{cases}1,&\text{if }D\phi(s_{t}^{T})>0.5,\\ -1,&\text{if }D\phi(s_{t}^{T})<0.5,\\ \text{any value in }[-1,1],&\text{if }D\phi(s_{t}^{T})=0.5.\end{cases}$ (16)
+Equation 16 implies the trainer assigns positive rewards to student state-action pairs that the discriminator assesses as more likely to be from the expert ($D_{\phi}(s^{T}_{t})>0.5$) and negative rewards to those unlikely to be from the expert ($D_{\phi}(s^{T}_{t})<0.5$).
+By this mechanism, the trainer’s policy optimization relies on the discriminator’s assessment to assign rewards that encourage expert-like behavior.
+All in all, the derivative of the trainer’s expected reward, Equation 13, with respect to its policy parameters is rewritten as:
+$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left((2D_{\phi}(s_{t}^{T})-1)a_{t}^{T}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
+The trainer adjusts $\pi_{T}$ to output high rewards when $D_{\phi}(s_{t}^{T})$ is high.
+Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to the expert behaviors, according to the discriminator.
+B.2 Lemma 2:
 The discriminator $D_{\phi}$, parameterized by $\phi$ will converge to a function that estimates the probability of a state-action pair being generated by the expert policy, when trained on samples generated by both a student policy $\pi^{\theta_{S}}$ and an expert policy $\pi_{E}$.
-Proof for Lemma 2: The training objective for the discriminator is framed as a binary classification problem over mini-batches from expert demonstrations and student-generated trajectories.
-The discriminator’s loss function $\mathcal{L}_{D}(\phi)$ is the binary cross-entropy loss, which for a mini-batch of size $n$ is defined as:
-$\mathcal{L}_{D}(\phi)=-\frac{1}{n}\sum_{i=1}^{n}\Big{[}y_{i}\log(D_{\phi}(s_{i},a_{i}))+(1-y_{i})\log(1-D_{\phi}(s_{i},a_{i}))\Big{]},$ (16)
-where $(s_{i},a_{i})$ are sampled state-action pairs from the combined replay buffer $\mathcal{D}=\mathcal{D}_{S}\cup\mathcal{D}_{E}$, with corresponding labels $y_{i}$ indicating whether the pair is from the expert $(y_{i}=1)$ or the student $(y_{i}=0)$.
-The stochastic gradient descent update rule for optimizing $\mathcal{L}_{D}(\phi)$ is then given by:
-$\phi\leftarrow\phi-\eta_{D}\nabla_{\phi}\mathcal{L}_{D}(\phi),$ (17)
-where $\eta_{D}$ is the learning rate for the discriminator.
-Through iterative updates, $D_{\phi}(s,a)$ will converge to $P(\pi_{E}|s,a)$, provided the minimization of $\mathcal{L}_{D}(\phi)$ progresses according to the theoretical foundations of stochastic gradient descent.
-The convergence relies on the assumption that the student’s policy $\pi^{\theta_{S}}$ and the expert policy $\pi_{E}$ induce stationary distributions of state-action pairs, such that the discriminator’s data source is consistently representing both policies over time.
-By minimizing $\mathcal{L}_{D}(\phi)$, we seek $\phi^{*}$ such that:
-$\phi^{*}=\arg\min_{\phi}\mathcal{L}_{D}(\phi).$ (18)
-Under typical conditions for convergence in stochastic gradient descent, the convergence to a local minimum or saddle point can be guaranteed.
-The discriminator’s ability to distinguish between student and expert pairs improves as $\mathcal{L}_{D}(\phi)$ is minimized, implying that $\lim_{n_{batch}\rightarrow\infty}D{\phi^{*}}(s,a)=P(\pi_{E}|s,a)$, where $n_{batch}$ is the number of batches.
-Appendix B Additional Experiments
-B.1 Noisy Expert Data
-(a) Expert traj.
-(b) RILe traj.
-(c) GAIL traj.
-(d) AIRL traj.
-(e) IQLearn traj.
-(f) RILe val.
-(g) GAIL val.
-(h) AIRL val.
-(i) IQLearn val.
-Figure 5: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death.
-(c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish.
-(b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal.
-Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively.
-The optimal path, derived from the reward of the trainer or discriminator, is highlighted with green lines.
-To demonstrate the advantage of using RL to learn the reward function in RILe, as opposed to deriving the reward directly from the discriminator in AIL and AIRL, we designed a 5x5 MiniGrid experiment.
-The grid consists of 4 lava tiles that immediately kill the agent if it steps in it, representing terminal conditions.
-The goal condition of the environment is reaching the green tile.
-The expert demonstrations are imperfect, depicting an expert that passes through a lava tile without being killed and still reaches the green goal tile.
-Using this data, we trained the adversarial approaches with a perfect discriminator, which provides a reward of $0.99$ if the visited state-action pair stems from the expert and $0.01$ otherwise.
-These values were chosen over 1 and 0 because both AIRL and GAIL use the logarithm of the discriminator output to calculate rewards.
-Results are presented in Fig.
-5.
-The value graphs (Fig.
-5e-g) are attained by computing the value of each grid cell $c_{i}$ as $\sum_{a\in A}D(c_{i},a)$ for AIRL and GAIL, and $\sum_{a\in A}\pi_{T}(c_{i},a)$ for RILe.
-Fig.
-5a shows the expert trajectory.
-GAIL (Fig. 5c), AIRL (Fig. 5d) and IQLearn (Fig. 5e) fail to reach the goal, as their agents either become stuck or are directed into lava.
-In contrast, RILe (Fig.
-5d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data.
-The difference in the value graphs between RILe and the baselines intuitively explains this outcome.
-In AIL and AIRL (Fig.
-5f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly.
-Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory.
-In contrast, RILe’s trainer agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting.
-By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig.
-5f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal.
-Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration.
-Since the student agent is guided by the trainer to also match the expert trajectory, it remains close to this path after passing the lava tiles.
-B.2 Robustness to Noise in the Expert Data
+Proof for Lemma 2: The discriminator’s objective is to distinguish between state-action pairs generated by the expert and those generated by the student.
+The training objective for the discriminator is framed as a binary classification problem over expert demonstrations and student-generated trajectories.
+The discriminator’s loss function $\mathcal{L}_{D}(\phi)$ is the binary cross-entropy loss, which is defined as:
+$L_{D}(\phi)=-\mathbb{E}_{(s,a)\sim p_{E}}[\text{log}(D_{\phi}(s,a))]-\mathbb{E}_{(s,a)\sim p_{\pi_{S}}}[\text{log}(1-D_{\phi}(s,a))].$ (18)
+where $p_{E}(s,a)$ is the state-action distribution of the expert policy, and $p_{\pi_{S}}(s,a)$ is the state-action distribution of the student agent.
+Considering that $x=(s,a)$, this loss can be rewritten as:
+[…]
+[…]
+As presented in [25], the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
+$D^{*}_{\phi}(x)$ $=\frac{p_{E}(x)}{p_{E}(x)+p_{\pi_{S}}(x)},$ (21)
+$D^{*}_{\phi}(s,a)$ $=\frac{p_{E}(s,a)}{p_{E}(s,a)+p_{\pi_{S}}(s,a)}.$ (22)
+This shows that the optimal discriminator estimates the probability that a state-action pair comes from the expert policy, normalized by the total probability from both expert and student policies.
+Appendix C Training Strategies
+The introduction of the trainer agent into the AIL framework introduces instabilities that can hinder the learning process. To address these challenges, we employ three strategies.
+Freezing the Trainer Agent Midway: Continuing to train the trainer agent throughout the entire process can lead to overfitting on minor fluctuations in the student’s behavior.
+This overfitting causes the trainer to assign inappropriate negative rewards, which diverts the student away from expert behavior—especially since the student agent may fail to interpret these subtle nuances correctly in the later stages of training.
+To prevent this, we freeze the trainer agent once its critic network within the actor-critic framework converges during the training process.
+We consider the trainer’s critic network to have converged when the change in the exponential moving average (with a smoothing factor of 0.99) of the critic output and its variance over a window of 50000 training iterations fall below a certain threshold.
+In all our experiments, this threshold is set to 0.1, which we found empirically after our hyperparameter search (see Appendix H).
+This threshold works for all settings where the reward is bounded between $-1$ and $1$, which is the case for all our experiments.
+Reducing the Trainer’s Target Network Update Frequency: We decrease the target network update frequency of the trainer agent to half that of the student agent.
+After our hyperparameter sweeps (see Appendix H), we empirically found that updating at half the student agent’s frequency works best.
+This adjustment aims to prevent overestimation bias in the trainer’s value function and to slow down its learning pace.
+By updating less frequently, the trainer provides more consistent and reliable reward signals.
+This steadier guidance helps the student agent better understand and adapt to the trainer’s rewards, facilitating more stable learning.
+Increasing the Student Agent’s Exploration: We increase the exploration rate of the student agent compared to standard AIL methods.
+We implement an epsilon-greedy strategy within the actor-critic framework, allowing the student to occasionally take random actions.
+This increased exploration enables the student to visit a wider range of state-action pairs.
+Consequently, the trainer agent receives diverse input, helping it learn a more effective reward function.
+This diversity is crucial for the trainer to observe the outcomes of various actions and to guide the student more effectively toward expert behavior.
+Appendix D Experimental Settings
+D.1 Evolving Reward Function
+We use single expert demonstration in this experiment.
+For RILe, we plot the reward function learned by the trainer.
+For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator.
+D.2 Reward Function Dynamics
+In this experiment, we select the student agent’s hyperparameters to be identical to those used in GAIL, ensuring that the only difference between the agents is the reward function.
+Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps in Appendix H.
+RFDC: We calculate the Wasserstein distance between reward distributions over consecutive 10,000-step training intervals, denoted as times $t$ and $t+10,000$.
+This metric quantifies how much the overall reward distribution shifts over time.
+Changes in reward distributions depend both on the reward function and the student policy updates.
+Since we use the same student agent with the same hyperparameters, higher RFDC values still indicate that the reward function is adapting more dynamically in response to the student’s learning progress.
+FS-RFDC: We compute the mean absolute deviation of rewards between consecutive 10,000-step training intervals for a fixed set of states derived from expert data.
+As the fixed set, we use all the states in the expert data.
+Since the states used for calculating rewards are fixed, changes in this value purely depend on the reward function updates.
+This metric assesses how the reward values for specific states change over time.
+CPR: We evaluate how changes in the reward function correlate with improvements in student performance.
+We store rewards from both the learned reward function and the environment-defined rewards in separate buffers.
+In other words, we collect samples from two reward functions: the learned reward function and the environment-defined reward function.
+The environment rewards consider the agent’s velocity and stability.
+Every 10,000 steps, we calculate the Pearson correlation between these rewards and empty the buffers.
+This metric evaluates whether increases in the learned rewards relate to performance enhancements.
+D.3 Motion-Capture Data Imitation for Robotic Continuous Control
+During training, we use 8 different random seeds and 8 distinct initial positions for the robot.
+The validation setting mirrors the training conditions: we sample initial positions from the same set of 8 possibilities and use the same random seeds.
+In this setting, the student agent selects actions deterministically, allowing us to assess its performance under familiar conditions.
+For the test setting, we evaluate the policy’s ability to generalize to new, unseen scenarios.
+We modify the initial positions of the robot by randomly initializing it in stable configurations not included in the fixed set used during training.
+Additionally, we use different random seeds from those in training, introducing new random variations that affect the environment’s dynamics during state transitions.
+This setup enables us to assess how well the learned policy performs when faced with novel initial conditions and environmental changes.
+D.4 Learning from Demonstrations
+Each method is trained using 25 expert trajectories provided in the IQ-Learn paper [7].
+We use single seed for the training, and after the training, run experiments with 10 different random seeds and report the mean and standard deviation of the results.
+D.5 Impact of Expert Data on Trainer-Student Dynamics
+In this experiment, both seeds and initial positions in the test setting are different from the training one, and we report values from the test setting.
+For every percentage of the expert-data in buffers, we continue trainings of both the trainer agent and the student agent of RILe.
+For instance, in 100% expert data in the trainer’s buffer case, both the student and the discriminator are trained normally using samples from the student agent.
+However, we didn’t include student’s state-action pairs to the trainer’s buffer, instead, we filled that buffer with a batch of expert data, and updated the trainer regularly using this modified buffer.
+Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student agent.
+However, student’s state-actions pairs are not included in the student’s buffer, and student agent is updated just by using expert state-action pairs, using rewards coming from the trainer agent for these expert pairs.
+Regarding the normalizations, we trained Behavioral Cloning (BC) and RILe across various data leakage levels, selecting the highest-scoring run (0% leakage RILe) as the baseline.
+Other scores and convergence steps are normalized by dividing by the score and convergence steps of the baseline (0% leakage RILe).
+For IQLearn, we used their reported numbers in their paper, as we couldn’t replicate their results with their code and hyperparameters.
+Appendix E Additional Experiments
+E.1 Robustness to Noise in the Expert Data
 To evaluate the robustness of RILe and baseline methods to noise in the expert data, we conducted experiments in the MuJoCo Humanoid-v2 environment.
@@
 IQ-Learn* 312 192 153 243 277
-B.3 Robustness of the Learned Reward Function
-We evaluated the robustness of the reward functions learned by RILe and AIRL [6] through an experiment similar to that conducted by [27].
+E.2 Robustness of the Learned Reward Function
+We evaluated the robustness of the reward functions learned by RILe and AIRL [8] through an experiment similar to that conducted by RHIRL [28].
 Initially, both methods were trained to learn reward functions in a noise-free MuJoCo Humanoid-v2 environment.
@@
 AIRL 5334 5005 4967
-B.4 Reward Curves
+Appendix F Extended MuJoCo Results
+We present MuJoCo results for the test setting, with standard errors, in Table 5.
+Table 5: Test results on four MuJoCo tasks with standard errors.
+RILe GAIL AIRL IQLearn DRAIL
+Humanoid-v2 5928 $\pm$ 188 5709 $\pm$ 63 5623 $\pm$ 252 327 $\pm$ 105 5755 $\pm$ 34
+Walker2d-v2 4435 $\pm$ 206 4906 $\pm$ 159 4823 $\pm$ 221 270 $\pm$ 43 4016 $\pm$ 127
+Hopper-v2 3417 $\pm$ 155 3361 $\pm$ 51 3014 $\pm$ 190 310 $\pm$ 47 1230 $\pm$ 73
+HalfCheetah-v2 5205 $\pm$ 31 4173 $\pm$ 94 3991 $\pm$ 126 755 $\pm$ 211 4133 $\pm$ 41
+F.1 Reward Curves
 We compare the reward curves of RILe, GAIL [5], AIRL [8], IQ-Learn [7], and AdapMen [29] in the MuJoCo Humanoid-v2 experiment.
 Since the task involves learning from expert trajectories, we combined AdapMen with an adversarial discriminator to enable training without an expert policy.
+Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen.
+AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
 As shown in the reward curves, despite RILe having multiple components, it is the most efficient method.
 This efficiency is achieved through the dynamic guidance of the trainer during training, which adapts the reward function to meet the student’s needs.
-Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen.
-AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
-Appendix C Compute Resources
+Appendix G Extended LocoMujoco Results
+We present LocoMujoco results for the validation setting and test setting, with standard errors, in Table 6 and 7, respectively.
+Table 6: Validation results on seven LocoMujoco tasks.
+RILe GAIL AIRL IQ BCO GAIfO DRAIL GAIL DRAIL RILe Expert
+Walk Atlas $895.4$$\pm 25$ $918.6$$\pm 133$ $356.0$$\pm 68$ $32.1$$\pm 4$ $28.7$$\pm 4$ $831.6$$\pm 41$ $741.3$$\pm 46$ $773.9$$\pm 13$ 1000
+Talos $884.7$$\pm 8$ $675.5$$\pm 105$ $103.4$$\pm 22$ $7.2$$\pm 2$ $19.9$$\pm 4$ $718.8$$\pm 16$ $963.7$$\pm 48$ $949.4$$\pm 54$ 1000
+UnitreeH1 $980.7$$\pm 15$ $965.1$$\pm 20$ $716.2$$\pm 124$ $12.5$$\pm 6$ $43.7$$\pm 8.4$ $586.6$$\pm 102$ $954.7$$\pm 20$ $973.5$$\pm 8$ 1000
+Humanoid $970.3$$\pm 101$ $216.2$$\pm 18$ $78.2$$\pm 6$ $6.8$$\pm 1$ $8.3$$\pm 1$ $345.7$$\pm 34$ $550.8$$\pm 148$ $595.3$$\pm 73$ 1000
+Carry Atlas $889.7$$\pm 44$ $974.2$$\pm 80$ $271.9$$\pm 30$ $39.5$$\pm 8$ $42.7$$\pm 9$ $306.2$$\pm 9$ $654.1$$\pm 109$ $344.1$$\pm 28$ 1000
+Talos $503.3$$\pm 72$ $338.5$$\pm 48$ $74.1$$\pm 8$ $11.7$$\pm 3$ $8.1$$\pm 1$ $444.5$$\pm 96$ $889.8$$\pm 163$ $874.3$$\pm 174$ 1000
+UnitreeH1 $850.6$$\pm 80$ $637.4$$\pm 90$ $140.9$$\pm 21$ $12.3$$\pm 2$ $30.2$$\pm 5$ $503.6$$\pm 55$ $620.8$$\pm 60$ $878.1$$\pm 46$ 1000
+Table 7: Test results on seven LocoMujoco tasks.
+RILe GAIL AIRL IQ BCO GAIfO DRAIL GAIL DRAIL RILe Expert
+Walk Atlas $870.6$$\pm 13$ $792.7$$\pm 105$ $300.5$$\pm 74$ $30.9$$\pm 10$ $21.0$$\pm 3$ $803.1$$\pm 68$ $834.4$$\pm 23$ 899.1 ±$\pm 17$ 1000
+Talos $842.5$$\pm 24$ $442.3$$\pm 76$ $102.1$$\pm 17$ $4.5$$\pm 3$ $11.9$$\pm 1$ $687.2$$\pm 44$ $787.7$$\pm 11$ $\mathbf{896.6}$$\pm 12$ 1000
+UnitreeH1 $966.2$$\pm 14$ $950.2$$\pm 13$ $568.1$$\pm 156$ $8.8$$\pm 3$ $34.8$$\pm 10$ $526.8$$\pm 72$ $940.8$$\pm 20$ $\mathbf{995.8}$$\pm 6$ 1000
+Humanoid $\mathbf{831.3}$$\pm 98$ $181.4$$\pm 24$ $80.1$$\pm 9$ $4.5$$\pm 2$ $3.5$$\pm 2$ $292.1$$\pm 25$ $814.6$$\pm 80$ $527.6$$\pm 39$ 1000
+Carry Atlas $\mathbf{850.8}$$\pm 62$ $669.3$$\pm 55$ $256.4$$\pm 47$ $36.8$$\pm 14$ $20.3$$\pm 1$ $402.9$$\pm 39$ $516.6$$\pm 60$ $317.1$$\pm 19$ 1000
+Talos $220.1$$\pm 88$ $186.3$$\pm 28$ $134.2$$\pm 18$ $10.5$$\pm 3$ $10.3$$\pm 2$ $212.5$$\pm 32$ $836.7$$\pm 160$ $\mathbf{840.5}$$\pm 133$ 1000
+UnitreeH1 $788.3$$\pm 71$ $634.6$$\pm 45$ $130.5$$\pm 22$ $14.4$$\pm 2$ $21.1$$\pm 6$ $504.5$$\pm 30$ $796.7$$\pm 131$ $\mathbf{909.5}$$\pm 9$ 1000
+Appendix H Hyperparameters
+We present hyperparameters in Table 8. For DRAIL, we replaced the discriminators with the implementation provided by DRAIL and adopted their hyperparameters for the HandRotate task.
+Our experiments revealed that RILe’s performance is particularly sensitive to certain hyperparameters. We highlight three key observations:
+  • RILe is more sensitive to the hyperparameters of the discriminator compared to other methods.
+Specifically, increasing the discriminator’s capacity or training speed, by using a larger network architecture or increasing the number of updates per iteration, adversely affects RILe’s performance.
+A powerful discriminator tends to overfit quickly to the expert data, resulting in high confidence when distinguishing between expert and student behaviors.
+This poses challenges for the trainer agent, as the discriminator’s feedback becomes less informative.
+  • The update frequency of the trainer agent’s target network influences the stability of the RILe framework.
+Lower update frequencies lead to improved stability.
+A slower-updating trainer provides more consistent reward signals, allowing the student agent to better adapt to the rewards.
+However, a lower update frequency slows down the learning process, as the trainer adapts more slowly to changes in the student’s behavior.
+Therefore, there is a trade-off between stability and learning speed that needs to be balanced.
+  • Enhancing the exploration rate of the student agent benefits RILe more than it does baseline methods.
+By encouraging the student to explore more, through strategies like higher entropy regularization or implementing an epsilon-greedy policy, the student visits a broader range of state-action pairs.
+This increased diversity provides the trainer agent with more varied data, enabling it to learn a more effective and robust reward function.
+The additional exploration helps the trainer to better capture the effects of different actions.
+Appendix I Compute Resources
 For the training of RILe and baselines, following computational sources are employed:
@@
   • 32GB Memory
-Appendix D Algorithm
+Table 8: Hyperparameter Sweeps and Best Hyperparameters for LocoMujoco and Humanoid Experiments
+Hyperparameters RILe GAIL AIRL IQ-Learn
+Discriminator Updates per Round 1, 2, 8 1, 2, 8 1, 2, 8 -
+Batch Size 32, 64, 128 32, 64, 128 32, 64, 128 -
+Buffer Size 8192, 16384, 1e5 8192, 16384, 1e5 8192, 16384, 1e5 -
+Network [512FC, 512FC] [256FC, 256FC] [64FC, 64FC] [512FC, 512FC] [256FC, 256FC] [64FC, 64FC] [512FC, 512FC] [256FC, 256FC] [64FC, 64FC] -
+Gradient Penalty 0.5, 1 0.5, 1 0.5, 1 -
+Learning Rate 3e-4, 1e-4, 3e-5, 1e-5 3e-4, 1e-4, 3e-5, 1e-5 3e-4, 1e-4, 3e-5, 1e-5 -
+Student Buffer Size 1e5, 1e6 1e5, 1e6 1e5, 1e6 1e5, 1e6
+Batch Size 32, 256 32, 256 32, 256 32, 256
+Network [256FC, 256FC] [256FC, 256FC] [256FC, 256FC] [256FC, 256FC]
+Activation Function ReLU, Tanh ReLU, Tanh ReLU, Tanh ReLU, Tanh
+Discount Factor ($\gamma$) 0.99, 0.97, 0.95 0.99, 0.97, 0.95 0.99, 0.97, 0.95 0.99, 0.97, 0.95
+Learning Rate 3e-4, 1e-4, 3e-5, 1e-5 3e-4, 1e-4, 3e-5, 1e-5 3e-4, 1e-4, 3e-5, 1e-5 3e-4, 1e-4, 3e-5, 1e-5
+Tau ($\tau$) 0.05, 0.01, 0.005 0.05, 0.01, 0.005 0.05, 0.01, 0.005 0.05, 0.01, 0.005
+Epsilon-greedy 0, 0.1, 0.2 0, 0.1, 0.2 0, 0.1, 0.2 0, 0.1, 0.2
+Entropy 0.2, 0.5, 1 0.2, 0.5, 1 0.2, 0.5, 1 0.05, 0.1, 0.2, 0.5, 1
+Trainer Buffer Size 8192, 16384, 1e5, 1e6 - - -
+Batch Size 32, 256 - - -
+Network [256FC, 256FC] [64FC, 64FC] - - -
+Activation Function ReLU, Tanh - - -
+Discount Factor ($\gamma$) 0.99, 0.97, 0.95 - - -
+Learning Rate 3e-4, 1e-4, 3e-5, 1e-5 - - -
+Tau ($\tau$) 0.05, 0.01, 0.005 - - -
+Entropy 0.2, 0.5, 1 - - -
+Freeze Threshold 1, 0.5, 0.1, 0.01, 0.001 - - -
+Appendix J Algorithm
 Algorithm 1 RILe Training Process
@@
 16:end for
+Algorithm 2 RILe Training Process with Off-policy RL
+1:Initialize student policy $\pi_{S}$, trainer policy $\pi_{T}$, and the discriminator $D$ with random weights.
+2:Initialize an empty replay buffers $B_{D}$, $B_{S}$, $B_{T}$ with different sizes
+3:for each iteration do
+4:     Sample trajectory $\tau_{S}$ using current student policy $\pi_{S}$
+5:     Store $\tau_{S}$ in replay buffers $B_{D}$, $B_{S}$, $B_{T}$
+6:     Sample a batch of transitions, $b_{S}$ from $B_{S}$
+7:     for each transition $(s,a)$ in $b_{S}$ do
+8:         Calculate student reward $R^{S}$ using trainer policy:
+$R^{S}\leftarrow\pi_{T}$ (26)
+9:         Update $\pi_{S}$ using calculated rewards
+10:     end for
+11:     Sample a batch of transitions $b_{D}$ from $B_{D}$
+12:     Train discriminator $D$ to classify student and expert transitions
+$\max_{D}E_{\pi_{S}}[\log(D(s,a))]+E_{\pi_{E}}[\log(1-D(s,a))]$ (27)
+13:     Sample a batch of transitions, $b_{T}$ from $B_{T}$
+14:     for each transition $(s,a)$ in $b_{T}$ do
+15:         Calculate trainer reward $R^{T}$ using discriminator:
+$R^{T}\leftarrow\upsilon(D(s,a))a^{T}$ (28)
+16:         Update $\pi_{T}$ using calculated rewards
+17:     end for
+18:end for

### v4

@@
 Inverse reinforcement learning (IRL) uncovers reward functions from expert demonstrations but relies on an iterative process that is often computationally expensive.
-Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons offer insufficient feedback for effective learning.
+Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons often offer insufficient feedback for effective learning.
 We introduce RILe (Reinforced Imitation Learning), a framework that combines the strengths of imitation learning and inverse reinforcement learning to learn a dense reward function efficiently and achieve strong performance in high-dimensional tasks.
@@
 Learning complex behaviors is critical for advancing artificially intelligent agents in fields such as robotics and strategic games.
-Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design, which is both time-consuming and error-prone.
+Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design.
+This is both time-consuming and error-prone.
 There are two ways to address the reward engineering problem.
-First, Inverse Reinforcement Learning (IRL) [1, 2] offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering.
+First, Inverse Reinforcement Learning (IRL) (Ng & Russell, 2000; Ziebart et al., 2008) offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering.
 IRL proceeds iteratively: it first trains a policy (the learning agent’s decision-making mechanism) using the current reward function, observes how well the agent’s behavior aligns with the expert’s, and then refines the reward function to better guide the policy toward expert-like behaviors.
 Repeating this process eventually yields a reward function capable of providing nuanced feedback at different stages of learning.
-However, this iterative procedure is computationally expensive [3], especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
+However, this iterative procedure is computationally expensive (Zheng et al., 2022), especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
 Second, Imitation learning (IL) bypasses explicit reward design by directly comparing learned behaviors to expert demonstrations via a comparison mechanism.
-Traditional IL approaches such as Behavioral Cloning (BC) [4] match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks.
-To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL [5], introduce a discriminator that operates as a comparison mechanism and judges how expert-like the learned behaviors are.
-However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements.
-Instead, they rely on a comparison metric, often a distance measure or binary classification, that merely checks whether the agent’s behavior is (or is not) similar to the expert.
-Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize.
-Consequently, both traditional IL and AIL struggle in high-dimensional environments [6, 7], where the agent needs more granular and adaptive feedback than these mechanisms provide.
-Adversarial Inverse Reinforcement Learning (AIRL) [8] attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator.
+Traditional IL approaches such as Behavioral Cloning (BC) (Bain & Sammut, 1995) match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks.
+To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL (Ho & Ermon, 2016), introduce a discriminator as a comparison mechanism that judges how expert-like the learned behaviors are.
+However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements, or boosts exploration.
+Instead, they rely on a distance measure or a binary classifier, that merely checks whether the agent’s behavior is (or is not) similar to the expert.
+Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize, or which actions to explore.
+Consequently, both traditional IL and AIL struggle in high-dimensional environments (Peng et al., 2018; Garg et al., 2021), where the agent needs more granular and adaptive feedback than these mechanisms provide.
+Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2018) attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator.
 However, AIRL tightly couples the reward function to the discriminator’s output, causing it to inherit AIL’s limitations in high-dimensional settings where more fine-grained guidance is needed.
-Real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog.
+In contrast, real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog.
 The teacher also refines how they teach as the student progresses.
-Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student’s learning progress more effectively.
-By contrast, existing approaches lack this continuous cooperative synergy.
+Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student more effectively.
+By contrast, existing approaches lack this cooperative synergy.
 Adversarial Imitation Learning (AIL) does update a discriminator alongside the policy, but the discriminator’s sole role is to distinguish expert-like behavior from non-expert behavior.
@@
 Meanwhile, IRL methods only refine the reward after the policy converges, missing the opportunity for continuous co-evolution throughout training.
-To address these issues, inspiring by these insights, we propose Reinforced Imitation Learning (RILe).
+To address these issues, inspired by these insights, we propose Reinforced Imitation Learning (RILe).
 RIle combines the adaptive reward benefits of IRL with the computational efficiency of AIL (Fig.
@@
 Specifically, the trainer queries a discriminator to measure how expert-like the student’s behavior is, then optimizes the reward function based on that feedback, without waiting for the policy to converge.
-RILe preserves nuanced reward shaping, while avoiding IRL’s heavy computational loop.
+RILe offers nuanced reward shaping, while avoiding IRL’s heavy computational loop.
 As a result, RILe is particularly effective in high-dimensional settings, where agents need fine-grained guidance at every stage of learning.
@@
 This avoids IRL’s repeated policy re-training and the purely discriminator-based rewards of AIL/AIRL.
-Reinforcement learning enables the trainer agent to explore multiple reward regimes, yielding a reward function that considers long-horizon effects.
+Unlike the competitive judge in AIL, using RL allows RILe’s trainer agent to explore reward strategies and learn adaptively from the student’s progress, establishing a cooperative dynamic that yields a reward function that considers long-horizon effects.
   2. 2.
@@
 Consequently, RILe enables accurate imitation of the expert performance in high-dimensional tasks.
-We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL, specifically GAIL [5] AIRL [8], GAIfO [9], BCO [10], IQ-Learn [7] and DRAIL [11].
-Our experiments spans four scenarios: (1) Empirically investigating how RILe’s reward-learning strategy differs from AIL and AIRL in a maze task, (2) Quantifying the adaptability of the learned reward function in a humanoid locomotion task, (3) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems, and (4) Analyzing the impact of explicit expert-data usage within RILe’s training process.
+We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL: GAIL (Ho & Ermon, 2016) AIRL (Fu et al., 2018), GAIfO (Torabi et al., 2018b), BCO (Torabi et al., 2018a), IQ-Learn (Garg et al., 2021) and DRAIL (Lai et al., 2024).
+Our experiments span six studies: (1) Empirically analyzing how RILe’s reward-learning differs from baselines, (2) Quantitatively analyzing the learned reward function in RILe, (3) Comparing different trainer-discriminator relationships in RILe, (4) Evaluating the noise robustness of RILe, (5) Analyzing the impact of using expert-data explicitly inside RILe, and (6) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems.
 Our results show RILe’s superior performance, particularly in high-dimensional environments, and highlight RILe’s ability to learn a dynamic reward function that effectively guides the student through multiple stages of training.
@@
 Imitation Learning
-Early work in IL introduced Behavioral Cloning (BC) [4], which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations.
-DAgger [12] refines BC by aggregating data over multiple iterations to mitigate compounding errors.
-GAIL [5] employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior.
-BCO [10] extends BC, and GAIfO [9] extends GAIL, both to handle state-only observation scenarios.
-DQfD [13] introduces a two-stage approach with pre-training, while ValueDice [14] aligns policy and expert distributions via a distribution-matching objective.
-More recently, DRAIL [11] leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
-Despite these advances, IL methods face challenges in high-dimensional environments [6, 7], where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance.
+Early work in IL introduced Behavioral Cloning (BC) (Bain & Sammut, 1995), which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations.
+DAgger (Ross et al., 2011) refines BC by aggregating data over multiple iterations to mitigate compounding errors.
+GAIL (Ho & Ermon, 2016) employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior.
+BCO (Torabi et al., 2018a) extends BC, and GAIfO (Torabi et al., 2018b) extends GAIL, both to handle state-only observation scenarios.
+DQfD (Hester et al., 2018) introduces a two-stage approach with pre-training, while ValueDice (Kostrikov et al., 2020) aligns policy and expert distributions via a distribution-matching objective.
+More recently, DRAIL (Lai et al., 2024) leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
+Despite these advances, IL methods face challenges in high-dimensional environments (Peng et al., 2018; Garg et al., 2021), where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance.
 RILe addresses these limitations through an adaptive trainer–student framework, where a learned reward function provides more nuanced guidance than standard IL comparison mechanisms.
 Inverse Reinforcement Learning
-Inverse Reinforcement Learning (IRL), introduced by [1], aims to uncover the expert’s intrinsic reward function from demonstrations.
-Major developments include Apprenticeship Learning [15], Maximum Entropy IRL [2], and adversarial variants like AIRL [8].
-IQ-Learn [7] reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability.
-More recent work focuses on unstructured data [16] and cross-embodiment transfer [17].
-Nonetheless, IRL methods struggle with computational inefficiency and limited scalability [18], particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly.
+Inverse Reinforcement Learning (IRL), introduced by Ng & Russell (2000), aims to uncover the expert’s intrinsic reward function from demonstrations.
+Major developments include Apprenticeship Learning (Abbeel & Ng, 2004), Maximum Entropy IRL (Ziebart et al., 2008), and adversarial variants like AIRL (Fu et al., 2018).
+IQ-Learn (Garg et al., 2021) reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability.
+More recent work focuses on unstructured data (Chen et al., 2021) and cross-embodiment transfer (Zakka et al., 2022).
+Nonetheless, IRL methods struggle with computational inefficiency and limited scalability (Arora & Doshi, 2021), particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly.
 RILe mitigates these challenges by jointly learning the policy and reward function in a single process, avoiding IRL’s iterative retraining loop and facilitating more efficient reward shaping for complex environments.
@@
 In this work, we consider $\gamma$-discounted infinite horizon settings.
-Following [5], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$.
+Following Ho & Ermon (2016), expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$.
 The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$.
-We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [19].
+We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ (Finn et al., 2016).
 Our work considers an imitation learning problem from expert trajectories, consisting of states $s$ and actions $a$.
@@
 3.2 Reinforcement Learning (RL)
-Reinforcement learning seeks to find an optimal policy, $\pi^{*}$.
-that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig.
+Reinforcement learning seeks an optimal policy, $\pi^{*}$.
+that maximizes the discounted cumulative reward from the reward function $R=R(s,a)$ (Fig.
 1-(a)).
-In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ [5, 20].
-The RL problem with a parameterized reward function and entropy regularization is defined as:
+In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ (Ho & Ermon, 2016; Bloem & Bambos, 2014).
+The RL problem with a parameterized reward function and entropy regularization is defined as
 $\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
@@
 GAIL trains both policy and the discriminator at the same time.
-AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green).
+AIRL implements a new structure on the discriminator, separating reward from environment dynamics by using two networks under the discriminator (see additional terms in green).
 (D) RILe: similar to IRL, learning a reward function from data.
@@
 Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$.
-Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [2] is defined as
+Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning (Ziebart et al., 2008) is defined as
 $\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(\mathbb{E}_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)
@@
 […]
-GAIL [5] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig.
+GAIL (Ho & Ermon, 2016) introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig.
 1-(c)).
@@
 In other words, the reward function that is maximized by the policy is defined as a similarity function, expressed as $R(s,a)=-\text{log}\ (D_{\phi}(s,a))$.
-AIRL [8] extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as:
+AIRL (Fu et al., 2018) extends AIL to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). AIRL structures the discriminator as:
 $D_{\phi,\psi}(s,a,s^{\prime})=\frac{\exp(f_{\phi}(s,a,s^{\prime}))}{\exp(f_{\phi}(s,a,s^{\prime}))+\pi(a|s)},$ (4)
@@
 To provide effective guidance, the discriminator must accurately identify whether a given state–action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$.
-GAIL [5] established the feasibility of such a discriminator (see Appendix B for details).
+GAIL (Ho & Ermon, 2016) established the feasibility of such a discriminator (see Appendix B for details).
 Trainer Agent
@@
 To address this, we define the trainer reward as:
-$R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$ (7)
+$R^{T}=e^{|\upsilon(D_{\phi}(s^{T}))-a^{T}|}$ (7)
 where $\upsilon(x)=2x-1$ scales the discriminator’s output, making it symmetric around zero.
 Including $a^{T}$ in the trainer’s reward ensures the trainer effectively learns from its own actions.
-Formally, the trainer’s objective is:
-$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
+Formally, we define the trainer’s objective as:
+$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[e^{|\upsilon(D_{\phi}(s^{T}))-a^{T}|}].$ (8)
 RILe
@@
 Both $\pi_{S}$ and $\pi_{T}$ can be trained via any single-agent RL method.
-The overall training algorithm is detailed in Appendix J.
+The overall training algorithm is detailed in Appendix I.
 The student agent aims to recover the optimal policy $\pi_{S}^{*}$:
@@
 Simultaneously, the trainer aims to recover $\pi_{T}^{*}$:
-$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}_{t}]\right].$ (10)
+$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[e^{|\upsilon(D_{\phi}(s_{t}^{T}))-a_{t}^{T}|}]\right].$ (10)
 By optimizing these objectives together, RILe efficiently learns both a reward function and a policy in high-dimensional settings where traditional AIL or IRL methods often struggle.
-Details on specific training strategies are provided in Appendix C.
+Achieving stable joint training requires specific techniques in RILe’s adaptive system, which are discussed in Appendix C.
 5 Experiments
-We evaluate the performance of RILe by addressing four key questions:
+We evaluate the performance of RILe in this section. We perform five ablation studies to analyze RILe:
   1. 1.
-    How does RILe’s reward-learning strategy differs from AIL and AIRL?
+    Reward Function Evaluation: We qualitatively analyze how RILe’s reward-learning strategy differs from baselines.
   2. 2.
-    How adaptive is RILe’s learned reward function?
+    Reward Function Dynamics: We quantitatively assess the reward function learned by RILe.
   3. 3.
-    How does RILe perform in both low- and high-dimensional continuous-control problem compared to AIL/AIRL/IRL?
+    Trainer-Discriminator Relation: We compare different trainer reward functions and investigate how the interaction between the trainer and the discriminator affects RILe’s performance
   4. 4.
-    How explicitly using expert-data within RILe’s training process affects the context-sensitive reward learning?
-Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC [4, 21], BCO [10]), adversarial imitation learning (GAIL [5], GAIfO [9] and DRAIL [11]), adversarial inverse reinforcement learning (AIRL [8]), and inverse reinforcement learning (IQ-Learn [7]).
-DRAIL [11] introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
-Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix H.
-5.1 Evolving Reward Function
-To answer the first question about RILe’s reward-learning strategy, we compare RILe’s performance with AI(R)L baselines in a maze setting.
-In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
+    Robustness to Noise and Covariance Shift: We investigate the robustness of RILe to different types of noise and covariate shift in the environment.
+  5. 5.
+    Explicit Usage of Expert Data: Analyzing the effect of using expert data explicitly inside RILe on the trainer-student dynamics.
+Then, we evaluate the performance of RILe in high-dimensional tasks with two experiments:
+  1. 1.
+    Learning from Expert Demonstrations: Learning from perfect expert demonstrations in a reinforcement learning benchmark.
+  2. 2.
+    Robotic Continuous Control from Motion Capture Data: Learning to walk and to carry a box from motion-capture data with four different embodiments with unique configurations.
 (a) RILe
@@
 Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards.
-Purple squares represent obstacles.
-RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
+Black squares represent obstacles.
+RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training and struggle to adapt.
+Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC (Bain & Sammut, 1995; Ross & Bagnell, 2010), BCO (Torabi et al., 2018a)), adversarial imitation learning (GAIL (Ho & Ermon, 2016), GAIfO (Torabi et al., 2018b) and DRAIL (Lai et al., 2024)), adversarial inverse reinforcement learning (AIRL (Fu et al., 2018)), and inverse reinforcement learning (IQ-Learn (Garg et al., 2021)).
+DRAIL (Lai et al., 2024) introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
+Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix G.
+5.1 Ablation Studies
+5.1.1 Reward Function Evaluation
+To qualitatively evaluate how RILe’s reward-learning strategy differs from AI(R)L baselines, we compare them in a maze environment.
+In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
 Figure 3 shows how each method’s learned reward function evolves during training.
@@
 As the student learns to reach the lower part of the maze, RILe shifts high-reward regions toward the top-left (second column), again encouraging incremental progress.
-The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around other key areas to keep the agent from getting stuck.
-Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance.
+The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around top-left areas to keep the agent from getting stuck.
+Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance by encouraging exploration.
 This dynamic reward adaptation gets important in higher-dimensional tasks as we show in Section 5.3.
-5.2 Reward Function Dynamics
-To address the second question on adaptability, we quantify how the reward function evolves and its relationship to the student’s performance.
-We compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
-We introduce three metrics (see Appendix D.2 for details): (1) RFDC (Reward Function Distribution Change): Wasserstein distance between reward distributions over consecutive training intervals, capturing overall shifts in reward space, (2) FS-RFDC (Fixed-State Reward Function Distribution Change): Mean absolute deviation of reward values at a fixed set of expert states over time, (3) CPR (Correlation between Performance and Reward): Pearson correlation between changes in the reward function and changes in the student’s performance.
 (a) RFDC
@@
 (c) Correlation between Performance and Reward (CPR): Pearson correlation between changes in the reward function and changes in the student’s performance.
-5.2.1 Adaptability of the Learned Reward Function
-We assess how adaptive the reward function learned by RILe is compared to that of AIL.
+5.1.2 Reward Function Dynamics
+To qualitatively evaluate how RILe’s reward function evolves during the training, we compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
+We introduce three metrics (see Appendix D.2 for details): (1) RFDC (Reward Function Distribution Change): Wasserstein distance between reward distributions over consecutive training intervals, capturing overall shifts in reward space, (2) FS-RFDC (Fixed-State Reward Function Distribution Change): Mean absolute deviation of reward values at a fixed set of expert states over time, (3) CPR (Correlation between Performance and Reward): Pearson correlation between changes in the reward function and changes in the student’s performance.
+Adaptability of the Learned Reward Function
+We compare the adaptability of the reward functions learned by RILe and AIL.
 Fig.
@@
 The advanced discriminator in DRAIL reduces the need for drastic reward function changes, yet RILe remains more adaptive than GAIL.
-Since the changing student policy indirectly affects the Fig.
-4(a), we also show changes in reward values for the fixed set of states in Fig.
+Since the changing student policy indirectly affects RFDC, we also show changes in reward values for the fixed set of states in Fig.
 4(b).
@@
 Higher adaptability of RILe ensures that the reward signal remains aligned with incremental performance improvements, enabling the student to receive more timely and effective guidance throughout training.
-5.2.2 Correlation between the Learned Reward and the Student Performance
+Correlation between the Learned Reward and the Student Performance
 We evaluate how changes in the reward function correlate with improvements in student performance.
@@
 While the discriminator’s reward signal effectively guides learning early on, its increasingly static nature at later stages fails to capture subtle performance improvements, leading to a negative correlation.
-5.3 Imitation in Continuous Control Tasks
-5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
-To answer the third question, we first test RILe in high-dimensional robotic locomotion tasks [22], where the agent must imitate motion-capture data for various robotic bodies.
+Figure 5: Trainer-Discriminator Relation: Comparison of different trainer reward functions, each defining a different relationship between the trainer’s action and the discriminator’s output.
+The student’s return curves on the left show how performance evolves, and the normalized final performance on the right presents a clear comparison between reward designs.
+Exponential naive converges faster but plateaus at a lower final reward, whereas exponential difference yields the highest performance.
+5.1.3 Trainer-Discriminator Relation
+We investigate how the interaction between the trainer agent and the discriminator affects RILe’s performance by comparing different trainer reward functions.
+Each reward function defines a different relationship between the trainer’s action $a^{T}$ and the discriminator’s output $D_{\phi}(s^{T})$.
+We consider following reward functions: (a) Difference ($R^{T}=-|\upsilon(D_{\phi}(s^{T}))-a^{T}|$), (b) Exponential Difference (default in RILe): $R^{T}=e^{-|\upsilon(D_{\phi}(s^{T}))-a^{T}|}$, (c) Multiplication ($R^{T}=\upsilon(D_{\phi}(s^{T}))a^{T}$)), (d) Naive ($R^{T}=D_{\phi}(s^{T})$)), (e) Exponential Naive ($R^{T}=e^{1-D_{\phi}(s^{T})}$)) and (f) Sigmoid (($R^{T}=D_{\phi}(s^{T})\sigma(a^{T})$))), where $\upsilon(x)=2x-1$ and $\sigma(x)=\frac{1}{1+e^{-x}}$.
+Figure 5 presents reward curves and normalized rewards, where all rewards are normalized according to the maximum mean achieved test reward.
+While the exponential naive reward function offers the fastest convergence, the exponential difference reward offers the best performance.
+Therefore, we use exponential difference reward as the default reward function in RILe.
+5.1.4 Robustness to Noise and Covariate Shift
+We evaluate RILe’s robustness to noisy expert demonstrations and environmental covariate shift.
+First, to evaluate the noise robustness of RILe In the MuJoCo Humanoid-v2 environment, we inject a zero-mean Gaussian noise (varying $\Sigma$) into either expert actions or states.
+We use GAIL (Ho & Ermon, 2016), AIRL (Fu et al., 2018), RIL-Co (Tangkaratt et al., 2021), IC-GAIL (Wu et al., 2019), and IQ-Learn (Garg et al., 2021) as baselines.
+Table 2 shows that RILe consistently outperforms baselines across all noise levels.
+Notably, it maintains high performance even under heavy noise ($\Sigma=0.5$).
+Second, inspired by Xu et al.
+(2022), we evaluate the stability of the reward functions learned by RILe and AIRL.
+First, we train both models in a clean environment.
+Then, we freeze the learned reward functions and train new student agents in environments where Gaussian noise is injected into their actions (covariate shift).
+Table 2 shows that the reward function learned by RILe demonstrates superior robustness to covariate shift, maintaining high performance even under increased noise levels.
+Table 1: Robustness to different noise levels in the expert demonstrations in MuJoCo Humanoid-v2 environment.
+RILe GAIL AIRL RIL-Co IC-GAIL IQ
+Noise Free $\Sigma=0$ 5928 5709 5623 576 610 327
+Action $\Sigma=0.2$ 5280 5275 4869 491 601 192
+$\Sigma=0.5$ 5154 902 4589 493 568 153
+State $\Sigma=0.2$ 5350 5147 4898 505 590 243
+$\Sigma=0.5$ 5205 917 4780 501 591 277
+Table 2: Robustness to covariate shifts in environment.
+RILe AIRL
+No Noise 5928 5623
+Mild
+$\Sigma=0.2$ 5201 5005
+High
+$\Sigma=0.5$ 5196 4967
+5.1.5 Explicit Usage of Expert Data inside RILe
+Figure 6: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps. Expert data usage speeds the training of RILe but reduce final performance.
+To assess the impact of using expert data explicitly inside RILe, we experiment in the MuJoCo Humanoid environment using a single expert trajectory from (Garg et al., 2021).
+We vary the proportion of expert data in the replay buffers from 0% to 100% (e.g., 25% indicates one quarter of each buffer is expert data; see Appendix D.5).
+Figure 6 shows that while more expert data in both the trainer’s and the student’s replay buffers accelerates RILe’s convergence, it reduces final performance.
+At 100% expert data, the student’s performance drops markedly.
+This indicates that excessive expert data hampers the trainer’s real-time adaptation, disrupting RILe’s context-sensitive reward customization.
+We compare RILe with IQ-Learn and BC, both of which rely heavily on expert data.
+Even with substantial amounts of expert data, RILe still performs better than baselines, indicating that RILe’s adaptive reward-shaping provides a crucial edge over those methods.
+Table 3: Test results on four MuJoCo tasks.
+RILe GAIL AIRL IQ
+Humanoid 5928 5709 5623 327
+Walker2d 4435 4906 4823 270
+Hopper 3417 3361 3014 310
+HalfCheetah 5205 4173 3991 755
+5.2 Learning from Expert Demonstrations
+We evaluate how RILe’s performance compares to baselines in a widely used learning from demonstration benchmark (Todorov et al., 2012; Brockman et al., 2016), where perfect expert state-action pairs are available.
+Table 3 shows RILe consistently achieves competitive or superior performance compared to AIL/AIRL/IRL methods.
+In particular, on high-dimensional tasks (e.g.
+Humanoid), RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function.
+5.3 Robotic Continuous Control from Motion Capture Data
+We test RILe in high-dimensional robotic locomotion tasks (Al-Hafez et al., 2023),for various robotic bodies.
+The agent must imitate motion-capture data, which is inherently noisy and only consists states without action information.
 This benchmark is especially demanding due to its complexity and dimensionality.
@@
 Overall, these results underscore that RILe outperforms the AIL/IL/IRL baselines, and benefits even more from the enhancements provided by the DRAIL variants, achieving performance levels close to the expert.
-The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role, suggesting avenues for future research.
+The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role.
 This overall performance aligns with our claim that an adaptive reward function is crucial for mastering complex, high-dimensional behaviors
-Table 2: Test results on four MuJoCo tasks.
-RILe GAIL AIRL IQ
-Humanoid 5928 5709 5623 327
-Walker2d 4435 4906 4823 270
-Hopper 3417 3361 3014 310
-HalfCheetah 5205 4173 3991 755
-Figure 5: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps, respectively. Expert data usage speeds the training of RILe but reduces final performance.
-5.3.2 Learning from Demonstrations
-We further evaluate RILe’s performance on four standard MuJoCo tasks [23, 24], Humanoid, Walker2d, Hopper, and HalfCheetah, with baseline comparisons widely reported in prior work.
-Table 2 summarizes test results.
-Overall, these results demonstrate that RILe consistently achieves competitive or superior performance compared to established AIL/AIRL/IRL methods.
-In particular, on high-dimensional tasks such as Humanoid, RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function in handling complex continuous control challenges.
-5.4 Explicit Usage of Expert Data
-To answer the fourth question, we use the Humanoid environment in MuJoCo [23, 24] with a single expert trajectory from [7].
-We vary the proportion of expert data in the replay buffers from 0% to 100%.
-For instance, 25% indicates a quarter of each buffer is expert data and 75% is agent data (see Appendix D.5 for details).
-Figure 5 illustrates how changing the proportion of expert data in both the trainer’s and the student’s replay buffers influences RILe’s training.
-Introducing more expert data accelerates RILe’s convergence (needs fewer training steps), but reduce the final performance.
-In the extreme case of 100% expert data, the student’s performance drops substantially.
-This indicates that too much expert data hamper the trainer’s ability to adapt to the student’s real-time needs, disrupting RILe’s context-sensitive reward customization.
-We also compare RILe’s results with IQ-Learn and BC, both of which rely heavily on expert data.
-Even when RILe is trained with substantial amounts of expert data, it still generally performs better, indicating that RILe’s adaptive reward-shaping provides a crucial edge over supervised methods.
 6 Discussion
 As our experiments demonstrate, RILe consistently outperforms baseline models across various tasks, thanks to its adaptive learning approach, where the trainer agent continuously adjusts the reward based on the student’s current learning stage.
-Our maze experiments provide an interpretable example of how the trainer agent tailors its rewards.
+Our maze experiment illustrates how the trainer agent tailors its rewards.
 By encouraging actions that might seem suboptimal for immediate imitation but advantageous for long-term learning, RILe establishes a curriculum that ultimately boosts performance.
-This adaptive strategy helps RILe to achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
-Nonetheless, policy stability remains challenging when the reward function is constantly evolving.
+This adaptive strategy helps RILe achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
+Nonetheless, policy stability remains challenging with dynamically evolving rewards.
 Freezing the trainer (see Appendix C) stabilizes learning but halts further adaptation, and the discriminator itself tends to overfit quickly.
-Future work could explore fully cooperative multi-agent RL techniques to allow ongoing adaptation, investigate ways to bound or regulate trainer updates, and consider discriminator-less formulations for reward learning.
+Future work may explore cooperative multi-agent RL to support continual adaptation, and consider discriminator-less formulations for reward learning.
 Despite these challenges, RILe shows that cooperatively learning the policy and the reward function can offer significant advantages over static or iteratively updated methods.
@@
 The trainer’s reward at timestep $t$ is:
-$r_{t}^{T}=\upsilon(D_{\phi}(s_{t}^{T}))a_{t}^{T}$ (11)
+$r_{t}^{T}=e^{-|\upsilon(D_{\phi}(s_{t}^{T}))-a_{t}^{T}|}$ (11)
 where $s_{t}^{T}=(s^{S}_{t},a^{S}_{t})$ is the trainer’s observation, $D_{\phi}(s_{t}^{T})$ is the discrimantor output that estimates the likelihood that $s_{t}^{T}$ comes from the expert data, and $\upsilon(D)=2D-1$ is a scaling function that maps discriminator’s output to the range $[-1,1]$.
@@
 The trainer aims to maximize $Q_{T}(s^{T}_{t},a^{T}_{t})$ to satisfy Equation 13.
-Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ has the same sign as $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
-$\alpha^{*T}_{t}=\begin{cases}1,&\text{if }D\phi(s_{t}^{T})>0.5,\\ -1,&\text{if }D\phi(s_{t}^{T})<0.5,\\ \text{any value in }[-1,1],&\text{if }D\phi(s_{t}^{T})=0.5.\end{cases}$ (16)
-Equation 16 implies the trainer assigns positive rewards to student state-action pairs that the discriminator assesses as more likely to be from the expert ($D_{\phi}(s^{T}_{t})>0.5$) and negative rewards to those unlikely to be from the expert ($D_{\phi}(s^{T}_{t})<0.5$).
+Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ matches $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
+$\alpha^{*T}_{t}=\upsilon(D_{\phi}(s^{T}_{t}))=2D_{\phi}(s^{T}_{t})-1$ (16)
+Equation 16 implies the trainer learns to match the discriminator’s scaled output.
 By this mechanism, the trainer’s policy optimization relies on the discriminator’s assessment to assign rewards that encourage expert-like behavior.
+Over time, this guides the student toward regions of the state-action space where $D_{\phi}(s^{T}_{t})\approx 1$, i.e., expert-like behavior.
 All in all, the derivative of the trainer’s expected reward, Equation 13, with respect to its policy parameters is rewritten as:
-$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left((2D_{\phi}(s_{t}^{T})-1)a_{t}^{T}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
+$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left(e^{-|\upsilon(D_{\phi}(s_{t}^{T}))-a_{t}^{T}|}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
 The trainer adjusts $\pi_{T}$ to output high rewards when $D_{\phi}(s_{t}^{T})$ is high.
-Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to the expert behaviors, according to the discriminator.
+Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to expert behaviors, according to the discriminator.
 B.2 Lemma 2:
@@
 […]
-As presented in [25], the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
+As presented in Goodfellow et al. (2014), the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
 $D^{*}_{\phi}(x)$ $=\frac{p_{E}(x)}{p_{E}(x)+p_{\pi_{S}}(x)},$ (21)
@@
 To prevent this, we freeze the trainer agent once its critic network within the actor-critic framework converges during the training process.
-We consider the trainer’s critic network to have converged when the change in the exponential moving average (with a smoothing factor of 0.99) of the critic output and its variance over a window of 50000 training iterations fall below a certain threshold.
-In all our experiments, this threshold is set to 0.1, which we found empirically after our hyperparameter search (see Appendix H).
-This threshold works for all settings where the reward is bounded between $-1$ and $1$, which is the case for all our experiments.
-Reducing the Trainer’s Target Network Update Frequency: We decrease the target network update frequency of the trainer agent to half that of the student agent.
-After our hyperparameter sweeps (see Appendix H), we empirically found that updating at half the student agent’s frequency works best.
-This adjustment aims to prevent overestimation bias in the trainer’s value function and to slow down its learning pace.
-By updating less frequently, the trainer provides more consistent and reliable reward signals.
-This steadier guidance helps the student agent better understand and adapt to the trainer’s rewards, facilitating more stable learning.
+Utilizing a Smaller Buffer for the Trainer Agent: We employ distinct replay buffer sizes for the student and trainer agents.
+We use larger buffer for the student compared to the trainer, as detailed in our hyperparameter configurations (see Appendix G).
+This strategy ensures the trainer primarily learns from the student’s recent interactions, allowing it to adapt its reward function more rapidly to the evolving student policy instead of optimizing based on potentially outdated historical data.
+This increased responsiveness provides more relevant, timely feedback to the student, which we found empirically contributes to more stable and effective co-adaptation within the RILe framework across different tasks.
 Increasing the Student Agent’s Exploration: We increase the exploration rate of the student agent compared to standard AIL methods.
@@
 In this experiment, we select the student agent’s hyperparameters to be identical to those used in GAIL, ensuring that the only difference between the agents is the reward function.
-Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps in Appendix H.
+Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps presented in Appendix G.
 RFDC: We calculate the Wasserstein distance between reward distributions over consecutive 10,000-step training intervals, denoted as times $t$ and $t+10,000$.
@@
 Additionally, we use different random seeds from those in training, introducing new random variations that affect the environment’s dynamics during state transitions.
-This setup enables us to assess how well the learned policy performs when faced with novel initial conditions and environmental changes.
+This setup enables us to assess how well the learned policy performs when faced with novel initial conditions.
 D.4 Learning from Demonstrations
-Each method is trained using 25 expert trajectories provided in the IQ-Learn paper [7].
+Each method is trained using 25 expert trajectories provided in the IQ-Learn paper Garg et al.
+(2021).
 We use single seed for the training, and after the training, run experiments with 10 different random seeds and report the mean and standard deviation of the results.
@@
 However, we didn’t include student’s state-action pairs to the trainer’s buffer, instead, we filled that buffer with a batch of expert data, and updated the trainer regularly using this modified buffer.
-Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student agent.
+Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student.
 However, student’s state-actions pairs are not included in the student’s buffer, and student agent is updated just by using expert state-action pairs, using rewards coming from the trainer agent for these expert pairs.
@@
 For IQLearn, we used their reported numbers in their paper, as we couldn’t replicate their results with their code and hyperparameters.
-Appendix E Additional Experiments
-E.1 Robustness to Noise in the Expert Data
-To evaluate the robustness of RILe and baseline methods to noise in the expert data, we conducted experiments in the MuJoCo Humanoid-v2 environment.
-Artificial noise sampled from a zero-mean Gaussian distribution with varying standard deviations ($\Sigma$) was added to a single expert trajectory, affecting either the actions or the states.
-The baselines used for comparison were GAIL [5], AIRL [8], RIL-Co [26], IC-GAIL [27], and IQ-Learn [7].
-As shown in Table 3, RILe consistently outperforms the baselines across different noise levels, demonstrating superior robustness even when a high amount of noise is present in the expert data ($\Sigma=0.5$).
-These results indicate that RILe is less sensitive to imperfections in the expert demonstrations compared to existing methods.
-Table 3: Test results in MuJoCo Humanoid-v2 environment, where artificial noise sampled from a zero-mean Gaussian distribution is added to a single expert trajectory.
-Results are aggregated over 20 different-seed environments.
-IQ-Learn* is trained using the official code and hyperparameters of the IQ-Learn algorithm.
-Noise-Free Action Noise State Noise
-$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$ $\Sigma=0.2$ $\Sigma=0.5$
-RILe 5681 5280 5154 5350 5205
-GAIL 5430 5275 902 5147 917
-AIRL 5276 4869 4589 4898 4780
-RIL-Co 576 491 493 505 501
-IC-GAIL 610 601 568 590 591
-IQ-Learn* 312 192 153 243 277
-E.2 Robustness of the Learned Reward Function
-We evaluated the robustness of the reward functions learned by RILe and AIRL [8] through an experiment similar to that conducted by RHIRL [28].
-Initially, both methods were trained to learn reward functions in a noise-free MuJoCo Humanoid-v2 environment.
-After training, these reward functions were frozen.
-Subsequently, new student agents were trained using these fixed reward functions in environments where Gaussian noise was added to the agents’ actions, with varying noise levels.
-Table 4 presents the results of this evaluation.
-The reward function learned by RILe demonstrates superior robustness to noise, maintaining high performance even under increased noise levels.
-In contrast, the performance of agents using the reward function learned by AIRL decreases more significantly as noise increases.
-These findings indicate that the reward function learned by RILe is more resilient to environmental noise, contributing to better agent performance in noisy conditions.
-Table 4: We test the robustness of learned reward functions.
-After training reward functions in a noise-free setting, reward functions are frozen, and used to train a new agent in a noisy environment, where Gaussian noise is added to agent’s actions in every step.
-No Noise Mild Noise High Noise
-$\Sigma=0$ $\Sigma=0.2$ $\Sigma=0.5$
-RILe 5748 5201 5196
-AIRL 5334 5005 4967
-Appendix F Extended MuJoCo Results
-We present MuJoCo results for the test setting, with standard errors, in Table 5.
-Table 5: Test results on four MuJoCo tasks with standard errors.
-RILe GAIL AIRL IQLearn DRAIL
-Humanoid-v2 5928 $\pm$ 188 5709 $\pm$ 63 5623 $\pm$ 252 327 $\pm$ 105 5755 $\pm$ 34
-Walker2d-v2 4435 $\pm$ 206 4906 $\pm$ 159 4823 $\pm$ 221 270 $\pm$ 43 4016 $\pm$ 127
-Hopper-v2 3417 $\pm$ 155 3361 $\pm$ 51 3014 $\pm$ 190 310 $\pm$ 47 1230 $\pm$ 73
-HalfCheetah-v2 5205 $\pm$ 31 4173 $\pm$ 94 3991 $\pm$ 126 755 $\pm$ 211 4133 $\pm$ 41
-F.1 Reward Curves
-We compare the reward curves of RILe, GAIL [5], AIRL [8], IQ-Learn [7], and AdapMen [29] in the MuJoCo Humanoid-v2 experiment.
-Since the task involves learning from expert trajectories, we combined AdapMen with an adversarial discriminator to enable training without an expert policy.
-Figure 6: Training reward curves for the MuJoCo Humanoid-v2 experiment comparing RILe, AIRL, GAIL, IQ-Learn*, and adapted AdapMen.
-AdapMen is combined with an adversarial discriminator to be able to train it without expert policy.
-As shown in the reward curves, despite RILe having multiple components, it is the most efficient method.
-This efficiency is achieved through the dynamic guidance of the trainer during training, which adapts the reward function to meet the student’s needs.
-Appendix G Extended LocoMujoco Results
-We present LocoMujoco results for the validation setting and test setting, with standard errors, in Table 6 and 7, respectively.
+Appendix E Extended LocoMujoco Results
+We present LocoMujoco results for the validation setting and test setting, with standard errors, in Table 5 and 6, respectively.
 Table 5: Validation results on seven LocoMujoco tasks.
@@
 UnitreeH1 $788.3$$\pm 71$ $634.6$$\pm 45$ $130.5$$\pm 22$ $14.4$$\pm 2$ $21.1$$\pm 6$ $504.5$$\pm 30$ $796.7$$\pm 131$ $\mathbf{909.5}$$\pm 9$ 1000
-Appendix H Hyperparameters
+Appendix F Extended MuJoCo Results
+We present MuJoCo results for the test setting, with standard errors, in Table 7.
+Table 7: Test results on four MuJoCo tasks with standard errors.
+RILe GAIL AIRL IQLearn DRAIL
+Humanoid-v2 5928 $\pm$ 188 5709 $\pm$ 63 5623 $\pm$ 252 327 $\pm$ 105 5755 $\pm$ 34
+Walker2d-v2 4435 $\pm$ 206 4906 $\pm$ 159 4823 $\pm$ 221 270 $\pm$ 43 4016 $\pm$ 127
+Hopper-v2 3417 $\pm$ 155 3361 $\pm$ 51 3014 $\pm$ 190 310 $\pm$ 47 1230 $\pm$ 73
+HalfCheetah-v2 5205 $\pm$ 31 4173 $\pm$ 94 3991 $\pm$ 126 755 $\pm$ 211 4133 $\pm$ 41
+Appendix G Hyperparameters
 We present hyperparameters in Table 8. For DRAIL, we replaced the discriminators with the implementation provided by DRAIL and adopted their hyperparameters for the HandRotate task.
@@
 This poses challenges for the trainer agent, as the discriminator’s feedback becomes less informative.
-  • The update frequency of the trainer agent’s target network influences the stability of the RILe framework.
-Lower update frequencies lead to improved stability.
-A slower-updating trainer provides more consistent reward signals, allowing the student agent to better adapt to the rewards.
-However, a lower update frequency slows down the learning process, as the trainer adapts more slowly to changes in the student’s behavior.
-Therefore, there is a trade-off between stability and learning speed that needs to be balanced.
+  • Employing distinct replay buffer sizes, particularly a smaller buffer for the trainer agent compared to the student agent, offers better stability.
+This encourages the trainer to learn primarily from recent student interactions, allowing its reward function to consider recent advancements in the student’s evolving policy, rather than optimizing based on potentially outdated historical data.
+This allows the trainer agent to be more responsive and provide more relevant, timely feedback.
   • Enhancing the exploration rate of the student agent benefits RILe more than it does baseline methods.
@@
 The additional exploration helps the trainer to better capture the effects of different actions.
-Appendix I Compute Resources
+Appendix H Compute Resources
 For the training of RILe and baselines, following computational sources are employed:
@@
 Freeze Threshold 1, 0.5, 0.1, 0.01, 0.001 - - -
-Appendix J Algorithm
+Appendix I Algorithm
 Algorithm 1 RILe Training Process